### v1

# Lee-Wick instability; pion condensate; strangelets; strange stars

Abstract
Researchers have long been interested in exotic states of matter. In the early 1970s, Migdal proposed existence of metastable or stable nuclei containing a pion condensate, and Bodmer posited existence of collapsed nuclei with a quark core. Lee and Wick proposed that compressed nuclear matter could undergo a phase transition into a scalar condensate state with effectively very light baryons. In the mid-1980s, Witten pointed out that strange quark matter may be an absolutely stable form of matter that can exist in the form of quark strangelets. These hypotheses stimulated large experimental efforts to detect these states in heavy-ion collisions and in terrestrial samples. Additionally, the cosmos is the arena for a broad search for traces of exotic matter in compact objects and other energetic phenomena. We review various approaches and arguments for a possibility of abnormal states of matter, including metastable and stable dense objects with condensates, as well as dilute nuclear objects. Among these, we discuss p-wave and s-wave pion condensations in nuclear systems and formation of a $\Delta$ resonance matter. Then we discuss a scalar mode condensation, which is triggered by the Pomeranchuk instability in dilute nuclear matter, and a clustering. We also mention recent ideas that a relativistic rotation and a dark-matter component may stabilize abnormal nuclear systems. Finally, we list some observational anomalies which cannot be yet appropriately explained by conventional physics.
keywords:
Lee-Wick instability; pion condensate; strangelets; strange stars
[…]
1 Introduction
A search for abnormal forms of matter accompanied the development of modern physics during the 20th century. In year 1931 Lev Landau wrote a paper Landau (1932), in which he derived an upper limit of star mass consisting of a ultra-relativistic gas $\sim 1.5\,M_{\odot}$ and hypothesized that heavier stars may possess regions in which the laws of quantum mechanics (and therefore of quantum statistics) are violated. In such “unheimliche Sterne” (Germ eerie stars) as they were dubbed by Landau, see Yakovlev et al. (2013), electroneutral mixture of protons and electrons, the only particles known by then, would be stable against particle annihilation. The discovery of neutron by James Chadwick in 1932 lead Dmitri Ivanenko to the idea Iwanenko (1932) that nuclei may consist of protons and neutrons instead of protons and electrons as thought before. In Heisenberg (1932a, b, 1933) Werner Heisenberg introduced the theory of an atomic nucleus consisting of nucleons, — strongly interacting particle with the putative isospin 1/2, which manifest themselves as protons and neutrons depending of the nucleon isospin projections. In 1934 Walter Baade and Fritz Zwicky Baade and Zwicky (1934) proposed that sufficiently massive stars end their life undergoing a transition (supernova event) into compact object consisting mainly of neutrons and called, therefore, neutron stars. More than three decades the neutron stars were considered as a theoretical exotica. Only 1967 Jocelyn Bell and Anthony Hewish reported Hewish et al. (1968) the first observation of a radio pulsar. By now neutron stars are not considered as exotica, and more than 3000 pulsars are observed, presently.
What is the ground state of the world around us has been one of the ‘big’ questions since the beginning of physics and then at the discussion of the role of the either. If there is some underlying symmetry behind the construction of our world, would it be possible to change the ground state? In condensed matter physics this concept was proposed by Landau in 1937 Landau (1937a, b) for the description of phase transitions. He pointed out that any phase transition may be characterized by a change in the value and/or the symmetry structure of some order parameter, which can be used as an effective degree of freedom in the thermodynamic potential. This order parameter can have a rather complicated relation to the fundamental degrees of freedom of the considered system. Generalization to the case of an order parameter interacting with the magnetic field was done by Vitaly Ginzburg and Lev Landau in 1950 Ginzburg and Landau (1950).
Addressing the question of a possible transformation of the ground state of our physical world it is natural to suggest a simplest choice that the order parameter should have quantum number of vacuum $0^{+}$ (spin and parity), i.e., correspond to some scalar field. There were no scalar particles known in the mid 60s. Nevertheless, Julian Schwinger tried to formulate a theory of strong interactions using the spin 0, spin 1, and spin 1/2 fields Schwinger (1957). The first attempts to build a model of interacting hadrons using a symmetry between vector and axial currents, which enter equally in weak interactions, were undertaken in works by John Polkinghorn, Murrey Gell-Mann and Maurice Lévy Adler and Bardeen (1969); Gell-Mann and Levy (1960) where the so-called $\sigma$-model, including scalar and pseudoscalar mesons coupled to nucleons, was formulated. The Lagrangian of this model contains the part symmetric with respect to local transformations keeping parity of the states and transformations mixing states with different parities, (the chiral invariant part) and the part, which is not invariant under the parity mixing transformations. Our physical ground state ($0^{+}$) is chosen as a result of the spontaneous symmetry breaking in the invariant part of the Lagrangian whereas the symmetry breaking part is treated as a rather weak perturbation. A dynamical mechanism of the nucleon mass generation through the mechanism of spontaneous breaking of the chiral symmetry was elaborated in independent works by Valentin Vaks and Anatoly Larkin Vaks and Larkin (1961) and Yoichiro Nambu Nambu (1960), applying superconductivity methods to a Heisenberg type four-fermion two-component Lagrangian.
We should certainly mention a role played by the spontaneous symmetry breaking in the unified electroweak theory of Abdul Salam, Sidney Glashow, and Steven Weinberg and the role played by the Higgs scalar boson in it Glashow (1959, 1961); Salam and Ward (1959, 1961); Weinberg (1967). The study of the properties of superdense matter described by unified gauge theories began in 1972 with the work of David Kirzhnits Kirzhnits (1972), who showed that the classical scalar field responsible for symmetry breaking disappears at a high enough temperature. In 1974 first grand unified theory was proposed by Howard Georgi and Sidney Glashow Georgi and Glashow (1974).
Now let us come to the subject of this review, discussion of yet considered, as exotica, possibilities of existence of various abnormal nuclear systems. Soon after prediction of quarks by Murrey Gell-Mann and George Zweig in 1964, Dmitry Ivanenko and Dmitry Kurdgelaidze suggested that a neutron star may transform in a quark star Ivanenko and Kurdgelaidze (1965), see also Itoh (1970). In 1971 Arnold Bodmer proposed that also atomic nuclei may undergo a transition in a superdense quark state Bodmer (1971).
In 1971 Arkady Migdal put forward the idea that ground state of atomic nuclei may contain a pseudoscalar condensate ($0^{-}$) of pions. He demonstrated occurrence of the pion condensation in a sufficiently deep potential well in scalar and electric fields forming deep potential wells for pions and suggested possibility of superdense stable or metastable nuclei glued by the pion condensate Migdal (1971). Three types of abnormal $\pi$ condensate nuclei: superdense, neutron and supercharged ones were discussed in Migdal (1974); Migdal et al. (1976), more details see in reviews Migdal (1978); Migdal et al. (1990).
In 1974 Tsung-Dao Lee and Gian-Carlo Wick suggested possible existence of a superdense state of nuclear matter characterized by a value of the scalar (e.g., $\sigma$) field different from its mean value in atomic nuclei Lee (1974); Lee and Wick (1974). This idea was reviewed in Lee (1975, 1976). The simplest variant of the model contains only two fields, the neutral scalar meson field and the isospinor nucleon field. As the $\pi$ condensate nuclei, the $\sigma$ abnormal nuclear state can be either stable or metastable. In the former case the normal nuclei after a while would undergo the phase transition to the new state. In first works the mean field approximation was used in spite of the assumed strong interaction of the scalar meson with the nucleons. To somehow justify this assumption the authors noted analogy between the abnormal state and the Bose–Einstein condensation one. Then Lee and Margulies argued Lee and Margulies (1975); Lee and Wick (1979) that quantum fluctuations may not change the statement about the existence of a super-dense state of nuclear matter.
At the workshop at Bear Mountain, to north from New York City, on “BeV/nucleon collisions of heavy ions” Lee presented his results on the possible vacuum reconfiguration Lee (1974) and noted Lederman and Weneser (1975) that “in order to produce the abnormal nuclear state, we must consider reactions in which (i) a large number of nucleons is involved and (ii) the nucleon density can be increased by a significant factor. One is, therefore, led to considerations of high energy collisions between heavy ions, say ${\rm U}+{\rm U}\to Ab+\dots$, where $Ab$ denotes the abnormal nuclear state.” The event at Bear Mountain was pivotal in the conception of heavy ion physics. 1974 there were the Synchrophasotron at the Joint Institute for Nuclear Research in Dubna and a high-energy heavy-ion accelerator, the Bevalac, being completed at the Lawrence Berkeley Laboratory in Berkeley. Ideas about possibility of the production of abnormal nuclei in heavy-ion collisions gave a push to active experimental investigations. One suggested also the look for collapsed quark nuclei Bodmer (1971) and superdense nuclei with pion condensate in heavy-ion collisions, see Refs. Aleshin et al. (1976); Galitskii (1976). A possibility of reaching high degree of compression of nuclear matter in shock waves that could help to from anomalously dense droplets in heavy-in collisions was investigated in Refs. Scheid et al. (1974); Sobel et al. (1975). Limits on the occurrence of abnormal nuclei in heavy-ion collisions and on the terrestrial concentration of abnormal nuclei were discussed Holt et al. (1976); Karnaukhov and Polikarpov (1977).
Cosmos was another domain to look for an anomalous state of matter. Yakov Zeldovich liked to claim that the Universe is a poor man’s accelerator: experiments don’t need to be funded, and all we have to do is to collect the experimental data and interpret them properly. Years later ideas of spontaneous symmetry breaking penetrated in cosmological models, cf. Linde (1990).
Probably in 1975, Migdal returned from a conference and told one of us (D.N.V.) that during his talk, in which he mentioned a possibility of the existence of anomalous supercharged nuclei with a pion condensate, Gell-Mann asked him about a role of electrons in such highly-charged systems. Migdal suggested to think about this problem. In works Migdal et al. (1976, 1977) the authors developed a relativistic semiclassical treatment of the Dirac equation in deep electric potentials and the many-particle approximation for electrons occupying levels of the lower continuum. The occupation of the levels by electrons and negative muons in the deep electric potential well was named “the fermion condensation”. The presence of the upper bound on the electric potential was discussed by Berndt Müller and Johann Rafelski in Müller and Rafelski (1975). Preparing this manuscript, we realized that Lee also discussed a role of electrons in abnormal nuclei. He wrote in Ref. Lee (1976) that “…the abnormal state can create $e^{+}e^{-}$ pairs. The $e^{+}$ will be sent to infinity, but a fair fraction of the $e^{-}$ will be kept within the abnormal nucleus. As $Z$ increases, the number of $e^{-}$ also increases. The interplay between the added Fermi energy of $e^{-}$ and the Coulomb energy may eventually bring the abnormal state to the point of instability when $A$ reaches $\approx 10^{4}$.” References Voskresensky et al. (1977); Voskresensky and Chernoutsan (1978); Voskresensky (1977) described the charged condensate pions and vacuum electrons in strong electric fields and suggested a possibility of existence of nuclei-stars of arbitrary size (up to the compact star size) glued by the charged $\pi$ condensate produced in the dense nuclear matter and the electrons produced from the vacuum.
In 1979 Mikhail Troitsky and Nikolai Chekunaev discussed in Troitsky and Chekunaev (1979) that at densities above the nuclear saturation density $n_{0}=0.16$ fm-3 the gain in the energy due to the pion condensation can exceed the energy loss associated with the transformation of nucleons into heavier baryon species like hyperons and $\Delta$ isobars. The similar idea was applied to quarks by Edward Witten in 1984. He conjectured in Witten (1984) that at increasing baryon density it becomes energetically favorable to occupy the $s$ quark Fermi sea. He argued that “strange matter” containing comparable numbers of up, down, and strange quarks may become absolutely stable. Edward Farhi and Robert Jaffe studied the strange matter with intermediate baryon number $10^{2}<A<10^{7}$ Farhi and Jaffe (1984), later Alvaro De Rujula and Sidney Glashow De Rujula and Glashow (1984); De Rujula (1985) suggested a possibility of existence of quark nuggets, larger size nuclearites and strange stars. The charge distribution in the quark nuggets and strange stars was considered in Alcock et al. (1986) in a similar manner as was done for supercharged nuclei in Refs. Migdal et al. (1977); Voskresensky et al. (1977); Voskresensky and Chernoutsan (1978). It was suggested that there may exist stars with hadronic shell and quark core –hybrid stars Shapiro and Teukolsky (1983); Bethe et al. (1987). First investigations were done ignoring possibility of color superconductivity. As a detailed review one may use Madsen (1999). Presence of the Cooper pairing in quark matter was known starting from works Barrois (1977); Bailin and Love (1984), where gaps were estimated as $\Delta\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}$ MeV, as for the nucleons. Then interest to possibility of quark stars was renewed after arguments were given in Rapp et al. (1998); Alford et al. (1998) in favor of large values of the quark-quark pairing gaps, $\Delta\sim 100$ MeV, see Alford et al. (2008) and references therein.
The axion quark nuggets (AQN) model was proposed in Refs. Zhitnitsky (2003, 2021); Zhitnitsky and Maroudas (2025). The latter model incorporates axion domain walls formed during the QCD phase transition. The domain walls act as stabilizing structures, alleviating the need for a first-order phase transition as in the original Witten model. A generic feature of the quark nuggets is the electrosphere. In case of antimatter axion quark nuggets the electrosphere of the $\sim 10^{-8}$ cm depth contains positrons. The axion quark nuggets behave as conventional dark matter components in the low density environment and become strongly interacting macroscopically large objects in the relatively high-density environment. The available energy due to matter-antimatter annihilation can be as high as $\sim 2$ GeV per baryon charge.
[…]
Although the initial studies Lee (1974); Lee and Wick (1974); Migdal (1974) discussed the possibility of the sigma or pion condensation occurring at densities both greater than and less than $n_{0}$, it soon became evident that the latter possibility was unrealistic. On the other hand various models of equation of state of the dilute nearly isospin-symmetric nuclear matter demonstrate the presence of region of the first-order phase transition and spinodal instability at nucleon densities $0.3n_{0}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}n\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}0.7n_{0}$ Röpke et al. (1982); Curtin et al. (1983); Bertsch and Siemens (1983); Schulz et al. (1983); Borderie et al. (2022). In this density region the compressibility of nuclear matter is negative and the scalar Landau-Migdal parameter of the Fermi liquid $f_{0}(n)$ is smaller than 1-1, that causes the Pomeranchuck instability. In Kolomeitsev and Voskresensky (2016) it was shown that the Pomeranchuck instability may lead to a condensation of an effective scalar field in a Fermi liquid. The presence of such condensate would remove instabilities in first and zeroth sounds. Owing to the condensate, a novel metastable state might exist at nuclear subsaturation density for $N\sim Z$, where $N$ is the number of neutrons and $Z$ is the number of protons. In Maslov and Voskresensky (2019) it was realized that the stability region remains also in case of significantly isospin-asymmetric matter.
It was found in Refs, Voskresensky (2024, 2025) that a rapidly rotating drop of nuclear matter, being not bound in the rest frame, may become bound in the rotation frame because of the formation of a charged pion condensate giant vortex. This observation invites looking for the nuclear drops with high angular momenta.
The existence of stable or metastable nuggets would have numerous important consequences in physics. Therefore, it would be worthwhile to continue theoretical and, especially, experimental investigations of such objects, although all of the above predictions are based on speculation about the possible behavior of dense hadron and quark matter that has not been tested against experimental data. In the body of this review we focus on some more specific details.
The paper has the following structure. In Sec. 2 we focus on the Lee-Wick model. Section 3 considers various aspects of the model of supercharged nuclei glued by light bosons. In Sec. 4 we discuss superheavy charged pion condensate nuclei, then $\pi^{0}$ ferromagnetism, and then we pay attention to subtleties associated with description of the $s$ wave pion-nucleon interaction and $s$ wave pion condensation. Section 5 discusses the $\Delta$ resonance matter. In Sec. 6 we focus on the hypothesis of strangelets. Section 7 presents the idea of existence of the dilute metastable nuclear state glued by a scalar condensate field. Section 8 discusses anomalies that defy conventional physics interpretations, part of which could be related to abnormal nuclear systems. The last Sec. 9 contains concluding remarks.
Throughout the text we use the system of units where $\hbar=c=1$.
2 The Lee-Wick model
Let us present the idea of Lee and Wick in a nut shell. Let us consider the Lagrangian density of nucleons interacting with a scalar field $\phi$:
$\mathcal{L}=\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-U(\phi)+\bar{\psi}(p\!\!\!/-m_{N}+g\phi)\psi\,$, (1)
where $m_{N}=939$ MeV is the nucleon mass, $g$ is the strong interaction coupling constant, and $U(\phi)$ is the scalar field potential. The scalar field $\phi$ will be treated on the mean-field level. Then the nucleon spectrum is $\sqrt{m_{N}^{*2}+p^{2}}$ with the effective nucleon mass $m_{N}^{*}=m_{N}-g\phi$. While in the vacuum $\phi=0$, Lee and Wick suggested that in the nucleon medium at some critical density the effective nucleon mass $m_{N}^{*}=m_{N}-g\phi$ can reach zero. It would occur for $\phi_{c}=m_{N}/g$. The energy density gain at this transition can be estimated as
$\Delta E(n)\approx U(\phi_{c})+\nu\frac{p_{\rm F}^{4}}{4\pi^{2}}-m_{N}\,n-\frac{3p_{\rm F}^{2}}{10m_{N}}n,$ (2)
where $p_{\rm F}=(3\pi^{2}n/\nu)^{1/3}$ is the nucleon Fermi momentum and $\nu=1$ for the neutron matter and $\nu=2$ for the isospin-symmetric nuclear matter. Then the critical density can be estimated as
$n_{c}=n_{c,0}\Big(1+\frac{3p_{{\rm F},c}}{4m_{N}}+\frac{9p_{{\rm F},c}^{2}}{20m_{N}^{2}}+O(p_{{\rm F},c}^{3}/m_{N}^{3})\Big),\quad n_{c,0}=\frac{U(\phi_{c})}{m_{N}}=\nu\frac{p^{3}_{{\rm F},c}}{3\pi^{2}}$ (3)
For a simplistic estimate we can take $U(\phi)=\frac{1}{2}m^{2}_{\phi}\phi^{2}$, where $m_{\phi}$ is the mass of the scalar meson. Then, neglecting the correction $O(p_{{\rm F},c}/m_{N})$ we find
$n_{c}\simeq n_{c,0}=\frac{m^{2}_{\phi}m_{N}}{2g^{2}}.$ (4)
Using typical values for the strong interaction constant $g\sim 10$ and the mass of the scalar meson, $m_{\phi}\sim m_{\sigma}\sim 700$ MeV, we obtain $n_{c}\sim 2n_{0}$. The energy gain per nucleon for $n>n_{c}$ and for $(n-n_{c})/n_{c}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}1$ can be estimated as $\mathcal{E}=\Delta E/n\simeq-(1-n_{c}/n)m_{N}$ . These rough estimates show to a possibility that the ordinary nuclei may be exist in a metastable state and after a while they may undergo the first-order phase transition to the superdense ground state. The energy of each nucleon inside the anomalous nucleus is $p$, and outside $\sqrt{m_{N}^{2}+p^{2}}$. Thus, there is a potential barrier holding the nucleons in the given abnormal state.
The main shortcoming of the above simplified argumentation is that we did not consider the equation of motion for the $\phi$ field following from the Lagrangian (1) or perform the minimization of the energy density with respect to $\phi$. To remove this shortcoming one has to consider potential $U(\phi)$, which has the spontaneous symmetry breaking pattern. Let us take the potential in the form
$U(\phi)=\frac{1}{2}\mu^{2}\phi^{2}-\frac{1}{\sqrt{2}}\mu\sqrt{\lambda}(1-\delta)\phi^{3}+\frac{1}{4}\lambda\phi^{4}\,,\quad\mu,\lambda>0\,$, (5)
which for $0<\delta<1-2\sqrt{2}/3\approx 0.057$ has two local minima: one at $\phi_{1}=0$ and another one at
$\phi_{2}=\frac{\phi_{0}}{4}(3(1-\delta)+\sqrt{9(1-\delta)^{2}-8})=\phi_{0}(1-3\delta+O(\delta)),\quad\mbox{where}\quad\phi_{0}=\sqrt{2}\frac{\mu}{\sqrt{\lambda}}.$ (6)
The global minimum of the potential is realized at $\phi=0$ that corresponds to the ground state. The second minimum is higher,
$U(\phi_{2})\approx\frac{1}{2}\lambda\phi_{0}^{4}(\delta+O(\delta^{2}))\,$, (7)
and the parameter $\delta$ is responsible for lifting the symmetry of two minima. In both minima the potential has the similar curvature $m_{\phi}^{2}=U^{\prime\prime}(\phi_{1})=\mu^{2}$ and $U^{\prime\prime}(\phi_{2})=\mu^{2}(1-12\delta+O(\delta^{2}))$. The potential is illustrated in Fig. 1.
Figure 1: Potential (5) for $\mu>0$, $\lambda>0$ and $0<\delta<1-2\sqrt{2}/3$.
The coupling of the scalar field to nucleons is defined as previously so that the nucleon mass becomes very small closely to the new minimum $\phi\sim\phi_{2}$,
$g=\frac{m_{N}}{\phi_{0}}=\frac{m_{N}}{\mu}\sqrt{\frac{\lambda}{2}}=\frac{m_{N}}{m_{\phi}}\sqrt{\frac{\lambda}{2}}\,.$ (8)
The critical density $n_{c,0}$ in (3) can be estimated now with the help of Eq. (7) as
$n_{c,0}=\frac{\lambda\phi_{0}^{4}}{2m_{N}}\delta=\frac{2m_{\phi}^{3}}{\lambda m_{N}}\delta=\frac{m_{\phi}^{2}m_{N}}{g^{2}}\delta\,$, (9)
where to get the second and third equalities we used Eqs. (6) and (8). In contrast to the more rough estimate (4) the critical density (9) is determined by the symmetry-breaking parameter $\delta$. Let us now consider the energy density of the system taking into account the nucleon kinetic energy
$E(\phi,n)=U(\phi)+\nu\intop_{0}^{p_{\rm F}}\frac{{\rm d}pp^{2}}{\pi^{2}}\sqrt{(m_{N}-g\phi)^{2}+p^{2}}$ (10)
and plot $\Delta E(\phi,n)=E(\phi,n)-E(0,n)$ as a function of $\phi$ for various values of the nucleon density. In Fig. 3 we present the results for the set of parameters
$\mu=m_{\phi}=800\,{\rm MeV}\,,\quad\lambda=10\,,\quad\delta=0.03\,.$ (11)
We see that with a density increase the relative height of the second minimum (at $\phi\neq 0$) decreases and at some density it becomes a global minimum. At this density the nucleon mass changes abruptly as shown in Fig. 3 by the solid line. The dashed line presents the result obtained with a value $\delta=0.02$. The transition density decreases, respectively.
In Fig. 4 we plot the binding energy per nucleon,
$\mathcal{E}(n)=\frac{1}{n}E(\min\phi(n),n)-m_{N}\,.$ (12)
We see, that for the parameter set (11), the vanishing nucleon mass leads to tremendous binding of the nuclear matter with a minimum realized at $\sim 12\,n_{0}$ with the binding energy of 340-340 MeV, see the solid line in Fig. 4. To reduce the binding one can take into account he nucleon-nucleon repulsion. To simulate it one can include a term quadratic in the nucleon density in expression (10),
$E(\phi,n)\to E(\phi,n)+\frac{1}{2}k\frac{n^{2}}{n_{0}},$ (13)
where $k$ is treated as a parameter. An increase of $k$ reduces binding, which disappears completely at the critical value, $k=98$ MeV, see the short-dashed line in Fig. 4. Reducing the value of $k$, we find that the minimum in the energy appears at $n\sim 5\,n_{0}$, and at the half critical value, $k=49$ MeV the minimum is located at the density $\sim 6.4\,n_{0}$ and has the depth of 136-136 MeV, see the dashed line in Fig. 4.
Figure 2: Difference in the energy density (10) due to the development of the scalar field as function of $\phi$ for various nucleon densities ($\nu=2$) calculated with the model parameters (11).
Figure 3: Effective nucleon mass as a function of the nucleon density, $\nu=2$. Solid line is calculated for the Lee-Wick model with parameters (11). Dashed line is drawn for $\delta=0.02$. Dash-dotted line corresponds to the RMF model (15) with parameters (16).
Figure 4: Binding energy per nucleon (12) for the nuclear matter ($\nu=2$) as a function of nucleon density. Difference in the energy density (10) due to the development of the scalar field is calculated for parameters (11). Solid line shows the result of the Lee-Wick model. Dashed lines include nucleon-nucleon interaction term (13). RMF curve shows RMF calculation with parameters (16).
The same year when Lee and Wick proposed their model, Walecka formulated a relativistic mean-field model (RMF) Walecka (1974); Chin and Walecka (1974). Here one should also mention previous works by Johnson and Teller, and by Dürr Johnson and Teller (1955); Duerr (1956), where the nuclear saturation arises due to the difference between scalar and vector nucleon densities. The Walecka model includes nucleons and mean fields of $\sigma$ and vector $\omega$ mesons allowing the $\sigma$ field dependence of the nucleon mass. Coupling constants of $\sigma$ and vector $\omega$ mesons with nucleons were fitted to describe binding energy and saturation density of nuclear matter. Incompressibility at nuclear saturation proved to be unrealistically high and effective nucleon mass low. To improve the fit Boguta included possible scalar field self-interactions depending on two extra fitting parameters Boguta and Bodmer (1977). Also $\rho$ meson mean field was included. These changes made the RMF models more flexible, see Serot and Walecka (1997). If in Eq. (10) we make replacements
$\phi\to\sigma\,,\quad\mu\to m_{\sigma}\,,\quad g\to g_{\sigma}\,,\quad\lambda\to cg_{\sigma}^{4}\,\quad\delta\to 1+\frac{\sqrt{2}}{3\sqrt{c}m_{\sigma}}bg_{\sigma}m_{N}\,$, (14)
and add the repulsive contribution due to the $\omega$ meson exchange with the mass $m_{\omega}$ and the coupling constant ($g_{\omega}$), we obtain the RMF energy density
$E(\sigma,n)=\frac{m_{\sigma}^{2}\sigma^{2}}{2}+\frac{b}{3}m_{N}(g_{\sigma}\sigma)^{3}+\frac{c}{4}(g_{\sigma}\sigma)^{4}+\nu\intop_{0}^{p_{\rm F}}\frac{{\rm d}pp^{2}}{\pi^{2}}\sqrt{(m_{N}-g_{\sigma}\sigma)^{2}+p^{2}}+\frac{g_{\omega}^{2}n^{2}}{2m_{\omega}^{2}}\,.$ (15)
The magnitude of the scalar field follows from the equation $\partial E(\sigma,n)/\partial\sigma=0$.
As example, taking the nucleon binding energy $\mathcal{E}_{0}=16$ MeV, compressibility modulus $K=275$ MeV and the effective nucleon mass $m_{N}^{*}(n_{0})=0.805$ at the saturation density $n_{0}$, as we did in Ref. Kolomeitsev and Voskresensky (2005), we obtain parameters
$\frac{g_{\sigma}^{2}m_{N}^{2}}{m_{\sigma}^{2}}=184.356\,,\quad\frac{g_{\omega}^{2}m_{N}^{2}}{m_{\omega}^{2}}=87.5996\,,\quad b=5.53871\times 10^{-3}\,,\quad c=2.29759\times 10^{-2}\,.$ (16)
The density dependence of the effective nucleon mass within the RMF model is shown by the dash-dotted line in Fig. 3. The mass decreases smoothly with the density increase without any jumps in difference to the Lee-Wick model. The binding energy per nucleon (12) for the nuclear matter ($\nu=2$) as a function of nucleon density is depicted in Fig. 4, solid line for Lee-Wick model. Long and short dashed lines include the nucleon-nucleon interaction term (13). The dash-dotted line shows the RMF calculation with parameters (16). The Lee-Wick abnormal state does not appear in the latter model.
Various modifications of the original Walecka model including $\rho$ mesons, a non-linear $\sigma$ field dependent non-linear potential, possible $\omega-\sigma$ interaction, etc, are widely and rather successfully used to describe equation of state of the hadron matter in compact stars, heavy-ion collisions and finite nuclei, cf. Serot and Walecka (1997); Geng et al. (2005); Voskresensky (1993) and references therein. In spite of all these goodies, for densities essentially above $n_{0}$ these models are phenomenological extrapolations. With parameters which were fitted to describe the nuclear system at $n=n_{0}$, at larger densities in these models the Lee-Wick abnormal state does not appear. However information about chiral symmetry is lost in these models, whereas the Lee and Wick model can be treated as a chiral symmetry based version of the RMF model. Pions are not included in both the Lee–Wick and RMF approaches, since they do not contribute on the mean-field level. Although not permitting to describe properties of the ordinary atomic nuclei and nuclear matter at $n\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}n_{0}$, the Lee-Wick model demonstrates principle possibility of existence of the new stable (or metastable) abnormal nuclear state at $n$ significantly above $n_{0}$. Note that one of ways to study the possibility of the Lee-Wick second minimum in $V(\sigma)$ within the relativistic mean-field approach to the equation of state of hadronic matter could be to employ within the latter approach the idea of a partial conservation of the chiral symmetry. It was realized in a series of works Kolomeitsev and Voskresensky (2005); Maslov et al. (2015, 2016); Kolomeitsev et al. (2017), where it was assumed that not only baryon but all hadron (baryon, $\omega$, $\rho$ and the $\sigma$ field) mass-terms decrease with increasing $\sigma$ field. Such assumption have permitted to successfully resolve several problems (associated with existence of very massive neutron stars, absence of the efficient direct Urca neutrino reaction in most neutron stars, so called hyperon and $\Delta$ isobar puzzles). For simplicity the parameters were chosen, which do not allow for the new minimum. Thus question about possibility of the Lee-Wick state within such kind of approach still remains open.
3 Supercharged nuclei, nuclearites and nuclei–stars
3.1 Idea of supercharged nuclei in a simplified treatment
For a moment, let us assume that negatively charged light scalar bosons (or pseudoscalar "light pions") exist with a mass of $m_{b}$. Could these bosons bind supercharged nuclei, as was suggested in Voskresensky (1977), see also review Migdal et al. (1990)? The energy of the nucleus with approximately constant density $n\simeq n_{0}$ at the proton number $Z\simeq A/2$ is given by
$\mathcal{E}$ $\approx-\mathcal{E}_{0}\,A+\mathcal{E}_{\rm surf}-\int n_{p}V{\rm d}^{3}x-\int\frac{(\nabla V)^{2}}{8\pi e^{2}}{\rm d}^{3}x$
$+\int\left(|\nabla\phi|^{2}-V^{2}|\phi|^{2}+m_{b}^{2}+{\textstyle\frac{1}{2}}\lambda|\phi|^{2})|\phi|^{2}\right){\rm d}^{3}x\,,\quad\mathcal{E}_{0}=16\,{\rm MeV}.$ (17)
Here the first term is the volume binding energy, the second term produces the surface energy, which is suppressed as $1/A^{1/3}$ compared to the first term and will be dropped when the nucleus atomic number $A$ is large as we assume. The proton density is $n_{p}\simeq n_{0}/2$. The third and fourth terms describe Coulomb interactions of protons and the electric field energy For ordinary atomic nuclei these terms produce the Coulomb energy $3Z^{2}e^{2}/5R$, where the nucleus radius is $R\simeq r_{0}A^{1/3}=(r_{0}/2^{1/3})Z^{1/3}$ and $r_{0}\simeq 1.2$ fm. The last term is the contribution of the classical charged boson field, the coupling of the pion-pion self-interaction $\lambda\sim 1$. With an increase in charge $Z\sim A/2$, the Coulomb energy repulsion always begins to exceed the volume energy attraction, so for the existence of extended systems with $Z\sim A/2>>1/e^{3}$ a charge must be ar least partially compensated. When $A$ and $Z$ grow, the electrical potential well deepens. When it becomes $V\leq-m_{b}$, the charged light bosons may appear in the reactions $n\to p+b^{-}$ and compensate the charge of the protons. As the result, in a nucleus with $A>>1/e^{3}$, the charge-neutral interior is formed Migdal et al. (1977) surrounded by a thin surface layer, where the strength of the electrical field is $E\sim 10^{2}m_{e}^{2}\sim 10^{18}$ V/cm with $m_{e}=0.5$ MeV being the electron mass.
The spatial structures of the electric potential $V(r)$ and the boson field $\phi(r)$ are governed by equations
$\Delta\phi+(V^{2}-m_{b}^{2}+\lambda|\phi|^{2})\phi=0\,$, (18)
$\Delta V=4\pi e^{2}(n_{p}+2V|\phi|^{2})\,.$ (19)
From these equations in case of a spherical nucleus with large $A$ we get uniform solutions $V=-m_{b}$ and $|\phi|^{2}=n_{p}/(2m_{b})$ valid in the interior, except for a boundary layer of the depth determined by the typical Debye screening length, being much shorter than the nucleus size. Here we put $\lambda\to 0$ to simplify consideration. Hence after the formation of the charged boson condensate the potential well stops growing keeping limiting depth of $V=-m_{b}$. Thus the volume part of the energy reads now
$\mathcal{E}=-\mathcal{E}_{0}A-VZ=-\mathcal{E}_{0}A+\frac{1}{2}m_{b}A\,.$ (20)
We conclude that if negatively charged scalar bosons with a mass of $m_{b}<2\mathcal{E}_{0}\simeq(30\mbox{--}32)$ MeV existed in nature, then stable supercharged nuclei and nuclearites of an arbitrary size (with $A$ up to $10^{56}$ and $n\sim n_{0}$) would also exist, being glued by strong and electromagnetic forces, see Refs. Voskresensky (1977); Migdal et al. (1990).
These exotic objects have been given various names: anomalous nuclei, abnormal nuclei, nuclear drops, and nuclei-stars. Similar self-bound objects made of quark were also called — nuggets, nuclearites, or strangelets. For larger-$A$ systems, gravity comes into the game, and we deal with compact stars with nucleon–light-boson (e.g., light pion) component. For objects build with the Witten’s strange quark matter, we will use further on the names like strangelets and strange stars.
We discussed abnormal nucleonic systems glued by negatively charged bosons but certainly, on equal footing, we could consider antinucleonic systems glued by positively charged bosons. In early Universe, the appearance of the latter anti-systems could contribute to a solution of the problem of the observable charge asymmetry of the Universe.
3.2 Dibaryons and neutral light bosons
The possibility of the existence of multiquark states was predicted by QCD-inspired models Jaffe (1977); Matveev and Sorba (1977). These works initiated a lot of experimental searches for six-quark states (dibaryons). Usually, one looked for dibaryons in the $NN$ channel. Such dibaryons would have decay widths from a few up to a hundred MeV. The obtained results were not very conclusive and the background determination was questionable. Enthusiasm has been revived in the field after experimental reports Tatischeff et al. (1997); Fil’kov et al. (2000, 2001). Reaction $pp\rightarrow p\pi^{+}\,X$ was studied in Tatischeff et al. (1997) and three narrow peaks (width $\sim 5$ MeV) in the missing mass spectrum were seen at $M_{X}=1004$ , $1044$ and $1094$ MeV with high statistical significance. The further study in Tatischeff et al. (2003) indicated the narrow baryonic structures at 1004, 1044, 1094, 1136, 1173, 1249, 1277, and 1384 MeV (and possibly 1339 MeV) in the missing mass $M_{X}$ and in the invariant mass $M_{p\pi^{+}}$. The reaction $pd\rightarrow ppX_{1}$ was studied in Fil’kov et al. (2000, 2001) and three peaks of the 5 MeV width were clearly observed in the missing mass $M_{pX_{1}}$ spectrum at $M_{pX_{1}}=1904\pm 2$ , $1926\pm 2$ and $1942\pm 2$ MeV. In the missing mass $M_{X_{1}}$ spectrum the peaks are located at $M_{X_{1}}=966\pm 2$ , $986\pm 2$ and $1002$ MeV. The peaks in the $pd$ reaction can be interpreted as the signatures of either dibaryons or nucleon resonances. Theoretical and experimental bounds on the properties of nucleon resonances were discussed in Ref. Azimov et al. (2003). The resonance states below the pion-nucleon threshold and their consequences for nuclear systems were discussed in Kolomeitsev and Voskresensky (2003).
Rather recently several very heavy compact stars were observed. The current highest precisely measured mass is $2.08(7)\,M_{\odot}$ for PSR J0740+6620 Cromartie et al. (2019); Fonseca et al. (2021). On the other hand appearance of hyperons and $\Delta$ resonances in neutron star interiors results in a decrease of the maximal neutron star mass. Nevertheless these so called the hyperon and $\Delta$ isobar puzzles can be resolved, see Kolomeitsev et al. (2017) and reference therein. Reference Kolomeitsev and Voskresensky (2003) demonstrated that in presence of light dibaryons the maximal possible mass of a neutron star would decrease below the observational limit. Appearance of rather heavy dibaryons and baryon resonances although results in a decrease of the maximal neutron star mass does not contradict to observations.
An interesting alternative interpretation of the tower of resonance states was given in Walcher (2001). It was noted that drawing all new states (both dibaryonic and nucleonic ones) on one energy scale, one obtains a tower of almost equidistant states. It was hypothesized the existence of a light pseudoscalar meson ($J^{P}=0^{-}$) with a mass of $m_{{\rm light}\,\pi}\simeq 21\pm 2.6$ MeV. The basic idea is that the tower of excited nucleonic states is formed by the nucleon in its ground state plus $1,2,3,\dots$ light pions as the quantum of excitation with the energy $m_{{\rm light}\pi}$. It was argued in Kolomeitsev and Voskresensky (2003) that the presence of light charged pions would allow for existence of abnormal nuclei ($A\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10^{3}$) and “nuclei stars” of arbitrary size, bound by strong and electromagnetic interactions.
Recent possible discovery is also worth mentioning: a light neutral boson (X17) with a mass of 17 MeV and a very narrow width (presumably less than $0.07$ eV) was found at the accelerator of ATOMKI. Three significant anomalies were observed in electron–positron pairs emitted in the 3H $(p,e^{-}e^{+})$ 4He, 7Li $(p,e^{-}e^{+})$ 8Be, and 11B $(p,e^{-}e^{+})^{12}$C reactions Krasznahorkay et al. (2016, 2021, 2022). If confirmed, the existence of this new particle could be important for particle physics and cosmology, see review Gustavino (2024). Reference Feng et al. (2016) assumed that X17 is the vector gauge boson coupled more strongly to neutrons than to protons. Reference Ellwanger and Moretti (2016) suggested another interpretation of the experimental results, assuming a light, pseudoscalar particle. In both cases the coupling of X17 with nucleons was assumed to be very weak.
On the other hand, independently of the value of the strength of the interaction it is worthwhile to ask, if existence of a light neutral boson does not contradict to the neutron star properties. Reference Kanakis-Pegios et al. (2024) assumed that X17 can mediate low-energy nucleon-nucleon interactions and attempted to find possible constraints on the hypothetical X17 boson from the analysis of the equation of state of neutron stars. It was shown that such a neutron star could have MRM-R curve rather similar to that for the ordinary neutron star.
Note that since $m_{X17}/m_{\pi}\sim 1/8$ one could expect presence of the nucleon–light boson polarization effects of the same order as for pions, if the coupling of the light boson with nucleons were $f_{XNN}\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}f_{\pi NN}/8$. In the latter case a possibility would exist for the Bose condensation of such a particles in the compact star matter (similarly to $\pi^{0}$ condensation) and feasibly for existence of metastable or stable nuggets and light-boson enriched nuclei-stars. Nevertheless we should repeat that Ellwanger and Moretti (2016) and Feng et al. (2016) estimated X17-nucleon polarization effects are tiny.
3.3 Rotating nuclear systems
The fastest known pulsar, PSR J1748-2446ad, makes one revolution in 1.4 ms (angular frequency $\Omega\approx 4.5\times 10^{3}$ Hz). Thus typical angular velocities of so far observed neutron stars are $\Omega R\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}{0.15c}$, where $c$ is the speed of light and $R\sim 10$ km is the neutron star radius. For a compact star rotating at the frequency close to the Kepler frequency $\Omega_{G}\sim\sqrt{GM/R^{3}}$, the rotation affects the equation of state resulting in an increase of the mass of the static star on the value $\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}(0.3-0.4)M_{\odot}$, cf. Farrell et al. (2024). When such a neutron star will undergo collapse to a black hole, its angular rotation frequency will increase due to conservation of the angular momentum up to $\Omega<1/r_{G}=1/(2MG)\sim 10^{5}$Hz for the compact star with the mass $M\sim M_{\odot}$, where $M_{\odot}\approx 2\times 10^{33}$ g is the solar mass, and $G$ is the gravitational constant. For primordial black holes with masses $M<10^{14}$ g, this estimate would give $\Omega\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}10^{24}$ Hz.
[…]
Condensation of pion excitations on a low-lying branch of the excitation spectrum in rectilinearly moving and rotating media was considered in Voskresensky (1993). Previously it was thought that a compact star becomes unstable provided the rotation frequency is enlarger than the Kepler value $\Omega>\Omega_{G}$. However Ref. Voskresensky (1993) argued that a part of the angular momentum can be transported to the forming condensate of Bose excitations, and as the result the angular frequency of the normal component may decrease below $\Omega_{G}$.
Magnetic fields in ordinary pulsars, like the Crab pulsar, reach values $\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}10^{13}$ G at their surfaces. At the surface of magnetars magnetic fields are higher, $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10^{15}$ G. In the interior the magnetic field might be even stronger (up to $\sim 10^{18}$ G) depending on the assumed mechanism of the formation of the magnetic field Voskresensky and Anisimov (1980). Also, magnetic fields $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}(10^{17}\mbox{--}10^{18})$ G, $h\sim h_{\rm VA}=Ze/R^{2}\sim H_{\pi}(Ze^{6})^{1/3}$ (where typical time of the collision process was estimated as $\Delta t\sim R/c$) may exist in non-central heavy ion collisions at collision energies $\sim$ GeV/A, as it was evaluated in Ref. Voskresensky and Anisimov (1980), $Z$ is the charge of the fireball, $H_{\pi}=m_{\pi}^{2}/e\simeq 3.5\times 10^{18}$ G.
In presence of a magnetic field and electrical potential well there exists the critical angular velocity, above which from the vacuum in the rotation frame there arises the pion condensate. In the frame rotating with the angular velocity $\vec{\Omega}\parallel z$ the solution for the charged condensate field has the form of the vortex, $e^{il\theta}\phi$, where $\theta$ is the angle in the cylindrical system and $l$ is the quantum number of the angular momentum. To be specific let further $\Omega,l>0$. Then equations for the boson field and the electric potential (18) and (19) are to be replaced by the following ones Voskresensky (2024, 2025),
$\Delta_{r}\phi+[(V-\Omega l)^{2}-m_{b}^{2}-(l/r-A_{\theta}(r))^{2}]\phi=0\,$, (21)
$\Delta_{r}V=4\pi e^{2}(n_{p}+2(V-\Omega l)|\phi|^{2})\,,\quad\Delta_{r}=\partial_{r}^{2}-\partial_{r}/r\,$, (22)
where $A_{\theta}(r)$ is the angular component of the vector potential of the magnetic field, and a $\phi$-field self-interaction is for simplicity neglected. Equations (21), (22) should be supported by equation for $\vec{A}$. However with the ansatz $A_{\theta}(r))=l/r$ we have $\mbox{curl}\vec{A}=0$ and equation for $\vec{A}$ is fulfilled identically.
Now in case of a nuclear system ($N\simeq Z$, $n\approx n_{0}$) of a large size solutions are $V=\Omega l-m_{b}\leq 0$, $|\phi|^{2}=n_{p}/(2m_{b})$ and instead of (20) we find
$\mathcal{E}=-\mathcal{E}_{0}A+\frac{1}{2}(m_{b}-\Omega l)A+\mathcal{E}_{kin}\,$, (23)
where ${\cal{E}}_{kin}$ is the kinetic energy of the rotation of the nucleon subsystem. The term $lA/2$ is smaller or equal to the initial angular momentum of the charged uniformly rotating nucleon sub-system $L\sim Am_{N}\Omega R^{2}$ (before appearance of the pion giant vortex, $l\gg 1$), $m_{N}$ is the mass of the nucleon. Thus $\Omega$ is determined by the initial value of the total angular momentum. Then $l\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}m_{N}\Omega R^{2}$. The critical value of $l$ is determined by the condition $\Omega l_{c}=m_{b}-2\mathcal{E}_{0}$. In the rotating frame the nuclearite remains bound till the system will not be decelerated by the weak radiation force. For a large system the latter time can be very long.
The rotation decreases the value of the effective pion mass, see (21), thereby favoring the pion condensation at a smaller critical density. This observation motivates a possibility of presence of the differential rotation characterizing by a more rapidly rotating charged pion superfluid (favoring a stronger condensate) and a more slowly rotating surrounding nucleon exterior. Notice that differentially rotating stars can support more mass than their uniformly rotating or non-rotating counterparts Morrison et al. (2004). In the case of massive remnant stars of binary merger events, differential rotation is one of the mechanisms that may provide an extra centrifugal support to stabilize the star well above the Tolman-Oppenheimer-Volkoff (uniform rotation) mass limit.
3.4 Anomalous supercharged nuclei consisting dark matter
In 1981 Cahn and Glashow Cahn and Glashow (1981) assumed existence of superheavy charged dark matter particles, cf. also De Rujula et al. (1990). In 2005 Glashow Glashow (2005) conjectured that more specifically the dark matter may consist of the U tera-quarks and O-UUU (charge 2-2) baryons, and in 2006 Khlopov Khlopov (2006) studied OOHe-hidden dark matter with $m_{O}\sim\!1$ TeV. Reference Gani et al. (2019) suggested existence of stable $O$-nuclearites of arbitrary size (up to $A\sim 10^{57}$), if there exist charged dark matter stable superheavy particles. In this case the proton charge of the nucleus is compensated by the heavy $O$-particles. Due to the large mass of the $O$ particles they yield a negligible contribution to the kinetic energy. Therefore, the ordinary isospin-symmetric matter at density $n\simeq n_{0}$ with embedded $O$-particles compensating the charge would be absolutely stable. Similarly to estimates (17), (20) performed above for supercharged nuclei, in case of the nuclear–$O$ matter one may present the energy as
${\cal{E}}\simeq\mathcal{E}_{0}\times A-\int d^{3}x(n_{p}-2n_{O})V\,.$ (24)
With $2n_{O}=n_{p}$ we get ${\cal{E}}\simeq\mathcal{E}_{0}A=-16A\,{\rm MeV}$.
At present there exist various anomalous events, some of which may have a relation to the discussed above scalar fields, $\pi^{-}$ condensate, quark or dark matter nuggets, see Zhitnitsky and Maroudas (2025) and references therein and a discussion in Sec. 8 below.
4 Superheavy nuclear objects
4.1 $\pi^{\pm}$ condensate nuclear systems
The phenomenon of a pion condensation was proposed and investigated in Refs. Migdal (1971, 1973); Scalapino (1972); Sawyer (1972), see also Migdal (1978); Migdal et al. (1990); Voskresensky (1993) and references therein. The pion-nucleon interaction is strongly attractive in the p-wave. Therefore, from the first works one expected that the pion condensate will be formed with a finite momentum $k\neq 0$ when the nucleon density exceeds some critical density $n_{c}^{p}$ Migdal (1973a, b); Baym (1973); Migdal et al. (1974); Campbell et al. (1975a, b); Baym et al. (1975); Dautry and Nyman (1979); Takatsuka et al. (1978). The $\pi N$ p-wave attraction is so strong that initially one anticipated that $n_{c}^{p}$ might be smaller than $n_{0}$. Soon, after performed analysis of the atomic nucleus data, it became clear that $n_{c}^{p}>n_{0}$ because of a stronger repulsive effect of the short-range nucleon-nucleon correlations, see  Migdal (1978); Migdal et al. (1990) and references therein.
Main contributions to the pion polarization operator come from the nucleon-nucleon hole and $\Delta$-nucleon hole terms both corrected by the short-range correlations, for more details see Migdal et al. (1990); Voskresensky (2023). The retarded pion propagator is
$D_{\pi}^{R}(\omega,\vec{k},n_{n},n_{p})=\frac{1}{\omega^{2}-m_{\pi}^{2}-k^{2}-\Pi^{R}(\omega,k,n_{n},n_{p})}\,$, (25)
where $\Pi$ is the pion polarization operator, $n_{n}$ and $n_{p}$ are the neutron and proton densities. For $\omega=\mu_{\pi}$:
$-[D_{\pi}^{R}(\mu_{\pi},\vec{k})]^{-1}=\widetilde{\omega}^{2}(\mu_{\pi},k^{2})=m^{2}_{\pi}+k^{2}+\Re\Pi^{R}(\mu_{\pi},k^{2})-\mu^{2}_{\pi}\,$, (26)
$\Re\Pi(\omega,k)$ is the real part of the pion polarization operator, the imaginary part $\Im\Pi(\mu_{\pi},k)=0$, $f_{\pi NN}\approx m_{\pi}^{-1}$ is the $\pi NN$ coupling constant, and $\Gamma(n)$ is the factor taking into account the $NN$ correlations, $\Gamma(n\to 0)\to 1$, $\mu_{\pi}$ is the pion chemical potential. In case of the isospin symmetric matter $\mu_{\pi^{\pm}}=0$ in accordance with reactions $p\leftrightarrow n+\pi^{+}$, $n\leftrightarrow p+\pi^{-}$; $\mu_{\pi^{0}}=0$ for the case of arbitrary isospin composition in accordance with reactions $N\leftrightarrow N+\pi^{0}$, $N=(n,p)$.
Spectra of pions in the isospin-symmetric and purely neutron matter were demonstrated in Migdal et al. (1974). In the latter case for $\pi^{\pm}$ there are the $\pi^{+}$ and $\pi^{-}$ branches, $\omega^{\pi^{+}}(k)$, $\omega^{\pi^{-}}(k)$ (corresponding to $\omega\to m_{\pi}$ for $n\to 0$) and the higher lying $\Delta$ branches. It proves to be that up to large densities $\omega^{\pi^{-}}(k)>\mu_{e}$ Migdal (1973a); Migdal et al. (1974) and the reaction $N+e\to N+\pi^{-}+\nu$ is forbidden. If pion were free, $\omega^{2}=m^{2}_{\pi}+k^{2}$, such a reaction were allowed for $\mu_{e}(n)>m_{\pi}$ resulting in efficient cooling of the neutron stars heaving such a densities in their interiors. For $n>n_{c}^{+}$ ($n_{c}^{+}<n_{0}$ according to evaluations of Migdal et al. (1974)) there appears extra branch with $\omega^{\pi^{+}}_{s}(k)<0$ corresponding to appearance of the so called $\pi^{+}_{s}$ condensate. Accordingly, for $n>n_{c}^{+}$ the reaction $p\to n+\pi^{+}$ becomes possible resulting in the production of $\pi^{+}$’s occupying this branch. For $n=n_{c}^{\pm}>n_{c}^{+}$ the branches $\omega^{\pi^{+}}_{s}(k)$ and $\omega^{\pi^{-}}(k)$ coincide at $k=k_{0}\neq 0$, where $\omega^{\pi^{+}}_{s}(k_{0})+\omega^{-}_{\pi}(k_{0})=0$. Thereby for $n>n_{c}^{\pm}$ the reactions $N\leftrightarrow N+\pi^{+}+\pi^{-}$ become energetically favorable and there appears $\pi^{\pm}$ condensate. Appearance of $\pi^{\pm}$ condensation for $n>n_{c}^{\pm}$ results in the efficient neutrino cooling processes, e.g. via $N+\pi_{cond}^{-}\to N+e+\bar{\nu}$ reaction.
The quantity
$\widetilde{\omega}^{2}_{0}=\min_{k}\{\widetilde{\omega}^{2}(\mu_{\pi},k)\}=\widetilde{\omega}^{2}(\mu_{\pi},k_{0})\,$, (27)
where the minimum is realized at a density dependent momentum $k_{0}(n)$, has the meaning of the squared effective pion gap, see Migdal et al. (1990) and references therein. We notice that the in-medium $\pi^{-}$ plays a role somewhat similar to the light negatively charged boson considered in the previous section, the boson squared mass, $m_{b}^{2}$, should be replaced by the squared effective pion gap.
At densities $n\sim n_{c}^{p}$ the pion gap behaves like $\widetilde{\omega}^{2}_{0}\sim\beta(n_{c}^{p}-n)$, with a constant $\beta(\mu_{\pi},k_{0})>0$ (at the neglect of the quantum fluctuations) and the change in the sign of the gap indicates the onset of instability with respect to the pion condensation with the momentum $k_{0}$. The critical point $n_{c}^{p}$ is the same for $\pi^{\pm},\pi^{0}$ in isospin symmetric matter. In the neutron star matter $n_{c}^{p}$ coincides with $n_{c}^{\pm}$ for charged pions and it is in general different for $\pi^{0}$.
In the isospin-symmetric matter the momentum dependence of the pion gap, $\widetilde{\omega}_{0}^{2}(k^{2})=\widetilde{\omega}^{2}(0,k^{2})$ develops a minimum at $k=k_{0}\neq 0$ at densities $n>n_{c}^{(1)}\simeq(0.5-0.7)\,n_{0}$, see Voskresensky (1993), and in its vicinity
$\widetilde{\omega}_{0}^{2}(k^{2})\simeq\widetilde{\omega}_{0}^{2}+\gamma\frac{(k^{2}-k_{0}^{2})^{2}}{4k_{0}^{2}}\,$,
where $\gamma\sim 1$, and $\widetilde{\omega}^{2}_{0}$ changes sign at $n=n_{c}^{p}$. Considering the pion condensation in a finite system for $n>n_{c}^{p}>n_{0}$, the spatial structure of the classical pion field is governed by the equation
$\widetilde{\omega}^{2}(\mu_{\pi},\hat{k})\phi+\lambda|\phi|^{2}\phi=0\,,\quad\hat{k}=-i\nabla\,$,
which replaces Eq. (18). Nuclear systems obeying the roton-like spectra have been extensively studied, see Voskresensky and Anisimov (1980); Migdal et al. (1990); Voskresensky (2023) and references therein. From a recent time some authors became to name such a spectrum the “moat” spectrum.
At large densities the baryon subsystem is essentially rearranged. In the approach exploiting the chiral symmetry  Campbell et al. (1975a, b); Baym et al. (1975) it was shown that instead of two different Fermi seas for protons and neutrons, a common Fermi sea for baryon quasiparticles, being superpositions of proton neutron and $\Delta$ isobar states mixed due to interactions with the developed pion condensate, is formed. At densities much larger than $n_{c}^{p}$ the squared amplitude of the pion condensate reaches its maximum value $f_{\pi}^{2}/2$ for $\pi^{\pm}$, where $f_{\pi}\approx 92$ MeV is the pion weak decay constant and the electric charge density of the baryon subsystem becomes equal to $n/2$. Taking into account these effects, and electron and muon populations, the energy density acquires the form Voskresensky et al. (1977)
${E}_{B+\pi}$ $={E}_{B}+{E}_{\pi}+{E}_{e}+{E}_{\mu}\,$,
${E}_{e,\mu}$ $\simeq\intop_{0}^{\sqrt{V^{2}-m_{e,\mu}^{2}}}\frac{{\rm d}pp^{2}}{\pi^{2}}\sqrt{m_{e,\mu}^{2}+p^{2}}\,\Theta(|V|-m_{e,\mu})\,$,
${E}_{\pi}$ $=\frac{V^{2}f^{2}_{\pi}}{2}+\frac{1}{3}(m^{*}_{\Delta}-m^{*}_{N})\,n-\frac{81}{50}f^{2}_{\pi NN}(1-g^{\prime})n^{2}\,$, (28)
where $E_{B}$ is the energy density of the baryonic subsystem, $g^{\prime}$ is the isospin Landau-Migdal parameter, and $m_{\Delta}^{*}$ is the effective mass of the $\Delta$ isobar (one may put for simplicity $m^{*}_{\Delta}-m^{*}_{N}\simeq m_{\Delta}-m_{N}=293$ MeV), $\Theta(x)$ is the step function. It is taken into account that in case of the developed condensate $\mu_{n}\simeq\mu_{p}$. For more details the reader may consult Refs. Voskresensky et al. (1977); Migdal et al. (1990). The electric charge of protons is screened by the condensate of negative pions and by electrons and muons produced via the weak $\beta$ processes and from the vacuum (at rather large values of ZZVoskresensky et al. (1977). The electric potential is found from the Poisson equation for a spherical system of radius $R$ and constant baryon density $n$,
$\Delta V$ $=4\pi e^{2}(n_{e}+n_{\mu}+n_{h})\,$,
$n_{h}$ $=\big({\textstyle\frac{1}{2}}n+f^{2}_{\pi}V\big)\Theta(R-r)\,$,
$n_{e,\mu}$ $=-\frac{1}{3\pi^{2}}\big(V^{2}-m_{e,\mu}^{2}\big)^{3/2}\Theta(|V|-m_{e,\mu})\,.$ (29)
Here $f^{2}_{\pi}V$ is the contribution of the developed $\pi^{-}$ condensate as it follows from the consideration within the $\sigma$ model and $n/2$ is the total baryon charge density in the presence of the developed condensate Baym et al. (1975), $n_{e}$ and $n_{\mu}$ are contributions of electrons and muons. Nullifying the r.h.s. of Eq. (29), similarly to that we did above for Eq. (19), demonstrates the charge neutrality condition in the interior of a large size system.
Figure 5: Binding energies per nucleon for the PNM (solid lines), the ISM (dashed lines), and for the beta-equilibrium matter ($pn\Delta$) with the developed $\pi^{-}$ condensate (dash-dotted lines). The latter lines are calculated for several values of the Landau-Migdal parameter $g^{\prime}$ characterising the strength of the short-range nucleon-nucleon correlations; the parameters are shown by labels on the lines. On the left panel, the nuclear part of the energy is taken from the FP equation of state Friedman and Pandharipande (1981). On the right panel we use the RMF model Cubero et al. (1988). Extra explanations are given in the paper body. The choice of parameters is such that the RMF equation of state supports the maximum neutron star mass of $2.09\,M_{\odot}$ Kolomeitsev and Voskresensky (2005).
In Fig. 5 by solid curves we show the binding energy per nucleon for the purely neutron matter (PNM). Other curves show calculations for the isospin symmetric matter (ISM) with and without the developed condensate. On left panel the nucleon part (dashed line) is calculated with a soft Friedman-Pandharipande (FP) equation of state Friedman and Pandharipande (1981) and on right panel with a stiffer RMF equation of state, see Eq. (15). In the RMF case parameters are tuned to describe the saturation density $n_{0}=0.160$ fm-3, the binding energy $\mathcal{E}_{0}=-16.0$ MeV, the compressibility modulus $K=273.6$ MeV, the effective nucleon mass $m_{N}^{*}(n_{0})=0.755$ MeV, and the value of the coefficient of the nuclear symmetry energy at saturation, $J=32.0$ MeV. With these parameters the maximum neutron star mass is $2.09\,M_{\odot}$, being larger $2\,M_{\odot}$ in agreement with the modern experimental data. The attraction strength of the $\pi^{\pm}$ condensate contribution, $E_{\pi}$ in Eq. (28), is controlled by the value of the $g^{\prime}$ Landau-Migdal parameter. The density dependence of this parameter is poorly known, therefore following Voskresensky et al. (1977); Migdal et al. (1990) we allow for its variation. So, the energies per nucleon of the baryon, $pn\Delta$, matter with the pion condensate (dash-dotted lines) are shown for several values of $g^{\prime}$. Note that various fits of the atomic nuclei data produce the value of $g^{\prime}(n_{0})\approx 0.7$. At this instance to avoid a possible misunderstanding we should mention that different authors use different normalizations of $g^{\prime}$ and distinct expressions for the pion polarization operator. The smaller $g^{\prime}(n)$ is the stronger is the pion condensate attraction for $n>n_{c}^{p}$. In Migdal et al. (1990) there were given some arguments for a possible decrease of $g^{\prime}(n)$ with increasing density. For the FP equation of state (left panel in Fig. 5) the energy of the pion condensate matter becomes negative in some interval of density for $g^{\prime}(n>n_{c}^{p})<0.482$. In this case there may exist anomalous superheavy pion condensate nuclei, nuclearites and nuclei-stars of arbitrary size bounded by nuclear forces rather than by gravity (for $10^{3}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}A\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}10^{56}$), as it was suggested in Ref. Voskresensky et al. (1977). The maximal net charge of superheavy nucleus with $A\gg 1/e^{3}$ observed at infinity behaves as $A^{1/3}$. In this case the extra charge is with necessity screened by electrons and muons occupying the vacuum shell. The pion condensate state becomes an absolute ground state on nuclear matter if $g^{\prime}<0.446$, then $\min_{n}\{{E}_{B+\pi}\}<-16$ MeV. For $0.482<g^{\prime}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}0.6$ there is a minimum in energy ${E}_{B+\pi}$ that would correspond to a metastable state. For the RMF nucleon equation of state (right panel in Fig. 5) the anomalous states may exist for $g^{\prime}<0.426$ and the absolute ground sate of matter appears for $g^{\prime}<0.381$. The metastable states are possible for $0.426<g^{\prime}<0.492$.
4.2 Superheavy $\pi^{0}$-condensate ferromagnetic nuclear systems
Possibility of neutron anomalous $\pi^{0}$ condensate nuclei was discussed in Migdal (1974); Migdal et al. (1976). Ferromagnetic behavior of $\pi^{0}$ condensate was studied within the $\sigma$ model first in Dautry and Nyman (1979). The $\sigma\pi^{0}$ running wave condensate in the neutron matter corresponds to a state, in which all neutron spins are aligned, presumably with some macroscopic domain structure. More recently there were found that extra contributions to the $p$ wave $\pi^{0}$ condensate term in the Lagrangian may arrive from the so-called Wess-Zumino-Witten axial anomaly term describing the anomalous interaction of the neutral pion field with the external electromagnetic field, and a related pion contribution to the baryon current, cf. Son and Stephanov (2008); Eto et al. (2013); Hashimoto (2015); Brauner and Yamamoto (2017); Huang et al. (2018).
Comparison of the ferromagnetic phase of the $\pi^{0}$ condensate appearing due to the axial anomaly term and the alternating layer structure of the pion condensate Takatsuka et al. (1978) demonstrated that the latter state is probably energetically more favorable Hashimoto (2015). These effects were recently studied in Ref. Voskresensky (2025).
4.3 About $s$ wave pion condensation
Pion off-mass-shell effects are of primary importance for description of the pion spectra and a possibility of the pion condensation. The most important contribution to the pion polarization operator is given by the $p$ wave pion-nucleon and the pion-$\Delta$-isobar interactions. For the $s$ wave part of the pion polarization operator already first works employed the Weinberg-Tomozawa expression, $\Pi^{s}_{\pi^{-}}=C(n_{n}-n_{p})\omega$, with $C\simeq 1/m^{2}_{\pi}$, which does not contribute in case of the ISM. Parametrization of the optical pion-nucleus potential Migdal (1978); Migdal et al. (1990); Voskresensky (1993) used the fully off-mass-shell pion-nucleon amplitude, which fulfills the current algebra theorems and the canonical PCAC condition (see the MSTV model in Voskresensky (2025)). Oppositely, Refs. Delorme et al. (1992); Ericson (1994); Kolomeitsev et al. (2003a, b) (see the KKW model in Voskresensky (2025)) used the on-mass-shell pion-nucleon amplitude, taking incoming and outgoing pion 4-momenta such that $q^{2}=q^{{}^{\prime}2}=m^{2}_{\pi}$, that does not allow to fulfill the so called Adler and Weinberg current algebra conditions. Subtleties associated with the current algebra theorems and field redefinitions were discussed in Voskresensky (2022). The $s$ wave pion-nucleon interactions in the linear sigma model, and in the Manohar-Georgi and Gasser-Sainio-Svarc models with finite number of terms in Lagrangians, as well as in a general phenomenological approach were reviewed.
With the MSTV model for the $s$ wave pion-nucleon interaction the $s$ wave pion condensation in the ISM does not occur at least up to very high densities. Oppositely, if the KKW model Delorme et al. (1992); Ericson (1994); Kolomeitsev et al. (2003a, b) were valid, the $s$ wave pion condensation in the ISM would be expected to occur already for $n>n_{c}^{s}\sim(1.4-2.5)n_{0}$, if the nucleon-nucleon correlations are not taken into account, resulting in the energy density gain
$E_{\pi}^{\rm KKW}(n)\simeq-\frac{m^{4}_{\pi}}{2\lambda}\left(\frac{n}{n_{c}^{s}}-1\right)^{2}\Theta(n-n_{c}^{s})\,.$ (30)
In Fig. 6 we plot the energy per particle of the ISM with and without the contribution of the $s$ wave pion condensation (30). We use the same RMF model as on the right panel of Fig. 5. The pion-pion self interaction parameter is chosen to be $\lambda=m_{\pi}^{2}/(2f_{\pi}^{2})$. The value of the softening of the equation of state essentially depends on the value of the critical density. As we see, for $n_{c}^{s}\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}2.5\,n_{0}$ the softening is weak, whereas for $n_{c}^{s}\leq 1.55\,n_{0}$ the effect is strong and there appears possibility of a metastable state with the $s$ wave pion condensate. Thereby one could expect to observe experimental manifestations of the $s$ wave pion condensate in heavy-ion collisions. We must, however, emphasize that the estimate of the value of the critical density $n_{c}^{s}$ remains strongly model-dependent even within the on-mass shell based model Delorme et al. (1992); Ericson (1994); Kolomeitsev et al. (2003a, b). For example, the $NN$ correlation effects, which cannot be constrained from the experiments at present, could shift the condensation critical point to higher values Voskresensky (2025).
Figure 6: Binding energy per nucleon for the ISM with and without the neutral $\pi^{\pm,0}$ condensate (solid and dash-dotted lines, respectively). The baryon part of the energy for ISM is calculated in the framework of the RMF model used on the right panel of Fig. 5. The dashed dotted curves are calculated within the KKW model for different values of the critical density $n_{c}^{\rm(s)}$ indicated by the line labels. In calculations we used $\lambda=m_{\pi}^{2}/(2f_{\pi}^{2})$.
5 $\Delta$ resonance matter
The idea of a stable or metastable baryon-resonance matter was introduced in Ref. Troitsky and Chekunaev (1979). The Fermi momentum of the nucleon is $p_{\rm F}\propto n^{1/3}/\nu^{1/3}$, the total Fermi energy of baryons grows as $E_{\rm F}\propto n^{2/3}\nu^{2/3}$ where $\nu$ is the degeneracy factor. When the nucleon Fermi energy, $E_{\rm F}$, reaches the difference between effective masses $m^{*}_{\Delta}-m^{*}_{N}$, it becomes energetically favorable to create $\Delta$ isobars. The appearance of any new degree of freedom results in a softening of the equation of state. The nucleon–nucleon-hole contribution to the pion polarization operator is $\Pi\propto-\nu p_{\rm F}k^{2}\propto\nu^{2/3}n^{1/3}$, see Voskresensky and Mishustin (1982). Therefore, the latter quantity effectively corresponds to the $\nu^{2}$ times higher density compared to the case of $\nu=1$. The extra $\Delta$$\Delta$-hole attractive term to the pion polarization operator leads to a further decrease of the critical density of the pion condensation. Moreover, let us recall that the spin-isospin degeneracy factor for $\Delta$ is $\nu=16$ compared to 4 for nucleons and that $f_{\pi N\Delta}\simeq 2f_{\pi NN}$.
$\Delta$ isobars are produced copiously in relativistic heavy-ion collisions forming a resonance matter Metag (1993). A possibility of density isomer states was discussed in Boguta (1982); Waldhauser et al. (1987); Cubero et al. (1988); Diakonov and Mirlin (1988) within RMF models and the chiral model of the nucleon.
[…]
6 Strangelets, hybrid and strange stars
6.1 Various models
Discussions of hypothetical strangelets and hybrid and quark stars began with the works Ivanenko and Kurdgelaidze (1965); Itoh (1970); Bodmer (1971); Witten (1984); Farhi and Jaffe (1984) and remain a hot topic till these days, see, e.g., Refs. Weber (2005); Farrell et al. (2024); Lugones and Grunfeld (2025); Di Clemente et al. (2025) and references therein. In Ref. Witten (1984) Witten estimated the energy per quark for a three-flavour system (strange quark matter) to be $\sim 90\%$ of the energy of a two-flavour system. This additional energy gain may compensate the energy loss due to the heavier strange quark mass leading, thereby, to stable quark matter.
The most commonly employed model of quark matter is the MIT bag model Chodos et al. (1974). By balancing the vacuum pressure outside the bag with the pressure of quarks inside it, one mimics phenomena of quark confinement and asymptotic freedom Farhi and Jaffe (1984). Besides the MIT-bag based approaches, the Nambu–Jona-Lasinio model is often used, see Ref. Klevansky (1992). The latter does not describe confinement of quarks and its preferred sets of parameters do not generally predict stable strange quark matter Buballa and Oertel (1999). The RMF approach proposed in Ref. Kaltenborn et al. (2017) gives a simple way to model confinement of quarks via introducing the particular density dependence of quark masses divergent for low baryon densities and a density-dependent screening effect. The screening is described analogously to an excluded volume effect for a hadronic matter model with density-dependent coupling constants. Effects of various pasta phases occurring in the quark-hadron phase transition Voskresensky et al. (2003) were then investigated in Ref. Maslov et al. (2019) for a set of RMF equations of state for both the hadron matter and the quark one. These approaches allow for hybrid stars but exclude stable quark star configurations. Some constraints may come from the lattice calculations as was discussed in Ivanov et al. (2005); Peshier et al. (2002); Schulze and Kampfer (2009). As a possible result, families of ordinary neutron stars and pure quark stars can be well separated. Within these models, the quark stars have low masses ($M\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}1M_{\odot}$), high densities ($n\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10n_{0}$), and small radii ($R\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}6$ km). Also, one uses the quark models with quark masses varying with the baryon number density, see Xia et al. (2014). Situation remains rather uncertain. Some works give arguments that even two-flavor quark matter can be stable or metastable, other argue against stable quark matter, e.g. see Holdom et al. (2018); Bai and Chen (2025) and references therein.
The quark objects are anticipated to be color superconductors. Various quark-paring patterns were investigated: color-flavor-locked phase, 2SC phase, spin-locking phase etc., see Alford et al. (2008); Farrell et al. (2024). Their effect on the equation of state is not significant even if the pairing gap is as large as $\Delta\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}100$ MeV. Owing a strong interaction a fluctuation region near critical point of the color superconductivity was estimated in Voskresensky (2004) as a broad and it is not excluded that some effects of the color superconducting fluctuations can be manifested in heavy-ion collisions.
Besides finite size strangelets and strange stars, also a branch of strange white dwarfs was also discussed Glendenning et al. (1995). Such dwarfs would have conventional nuclear crusts and strange quark matter cores. Possible candidates for strange dwarfs were suggested in Ref. Kurban et al. (2022). Potential implications for gravitational wave astronomy are highlighted in Ref. Perot et al. (2023).
Recently, a novel state of deconfined matter was proposed in Ref. Zhitnitsky and Maroudas (2025) using the model described in Ref. Zhitnitsky (2003). This proposed object consists of a core of quarks (antiquarks) in a color-superconducting state surrounded by an electrosphere of positrons (electrons), and enclosed by an axion domain wall. Large axion domain walls can form large bubbles filled with $u$, $d$ and $s$ quarks. Earlier, the idea that soliton-like configurations could serve as dark matter arose in Refs. Friedberg et al. (1976); Coleman (1985); Lee (1987); Lee and Pang (1989, 1992). Among other conditions, the possibility of existence of these objects is based on the feasible existence of a conserved charge associated with an ungauged unbroken continuous internal symmetry. In Ref. Coleman (1985) such objects were called $Q$ balls. Large $Q$ balls may form so-called boson stars. A review of these topics can be found in Shnir (2018); Nugaev and Shkerin (2020). Dark matter can be associated with a light boson that has a mass $m\sim 10-22$ eV and a de Broglie wavelength of $\sim 1$ kpc. This is often called the fuzzy dark matter. A soliton is surrounded by an envelope resembling a cold dark matter hallo Hui et al. (2017).
We should mention that recently, the mass and radius of the Type-I X-ray burster and accretion-powered millisecond pulsar XTE J1814-138 were inferred to be $M=(1.21\pm 0.05)M_{\odot}$ and $R=7.0\pm 0.4$ km (68.3% credibility interval), using the pulse profile modeling technique Kini et al. (2024). If were confirmed, the compact object of such a small radius cannot be described by conventional purely hadronic equations of state of the neutron stars. This could then be a candidate for the strange star, the pion condensate nucleus-star, the Lee-Wick $\sigma$ condensate compact object, or something else.
6.2 Cooling of strange stars and strangelets
The neutrino emissivity of quark matter was first considered by Iwamoto in Ref. Iwamoto (1982). In Ref. Pizzochero (1991) the surface temperature of the strange star was found to be lower than that of a neutron star of the same age (ordinary or with mesonic condensates). The effect of color superconductivity on the cooling of strange and hybrid stars has been studied in Refs. Blaschke et al. (2000, 2001); Grigorian et al. (2005) and subsequent works. Some models show that cooling can be so fast that strange stars may presently have temperatures smaller than the cosmic microwave background radiation Shaulov et al. (2022); Kupriyanova and Shaulov (2024).
Since we have mentioned the axion-based model of nuggets, we should also mention a possible axion contribution to cooling of neutron stars. The emission of axions by neutron stars was, first, discussed in Ref. Iwamoto (1984) without the inclusion of $NN$ interactions. Then in Ref. Voskresensky and Senatorov (1986) it was shown that axion emission should be strongly affected by in-medium baryon-baryon interactions similarly to the corresponding reaction of the neutrino production.
7 Dilute stable nuclear systems
7.1 Again scalar condensate
The analysis of two independent nuclear emulsion experiments exposed to beams of 16O and 56Fe ions with 2A GeV kinetic energy allowed for the interpretation of some events as the relatively rare occurrence of anomalous nuclear fragments interacting with an unexpectedly large cross section as described in Ref. Friedlander et al. (1983). These objects were called anomalons Bayman and Tang (1987). Although these results were not reproduced by subsequent measurements, uncertainty remains.
Below we discuss a possibility of existence of very dilute stable or metastable nuclear systems. We will consider the nuclear matter at zero temperature as a normal Fermi liquid. The quasiparticle part of the nucleon Green’s function close to the Fermi surface can be presented as
$G_{N}(\epsilon,\vec{p},n)=\frac{a}{\epsilon-\xi_{\vec{p}}+i\,0\,{\rm sign}\epsilon}\,,\quad\xi_{\vec{p}}=\frac{p^{2}-p_{\rm F}^{2}}{2\,m_{N}^{*}}\approx v_{\rm F}(p-p_{\rm F})\,,\quad v_{\rm F}=p_{\rm F}/m_{N}^{*}\,.$ (31)
The non-pole part of the Green’s function is hidden in the quasiparticle renormalization factor $a$, effective nucleon mass $m_{N}^{*}$ and the quaisparticle interaction amplitude in the particle-hole channel, $\Gamma_{0}^{\omega}=f_{0}/(a^{2}N)$, which we are interested here. The interaction is expressed here through the scalar Landau-Migdal parameter $f_{0}$ and the density of states at the Fermi surface $N=\nu m_{N}^{*}p_{\rm F}/\pi^{2}$. For ISM the degeneracy factor $\nu=2$. The parameter $f_{0}$ is related to the compressibility of the matter as follows
$K_{N}=n\frac{{\rm d}^{2}E_{N}}{{\rm d}n^{2}}=\frac{p_{\rm F}^{2}}{3m_{N}^{*}}(1+f_{0})\,$, (32)
where $E_{N}$ is the energy density of the nucleon system. Parameter $K_{N}$ is related to the standard compressibility modulus of the nuclear matter at saturation that we mentioned in Section 2 as $K=9K_{N}$.
The particle-hole scattering amplitude in the spin zero channel on the Fermi surface can be written throught a propagator of the scalar bosonic mode Kolomeitsev and Voskresensky (2016):
$T_{\rm ph,0}=-{\rm sgn}(f_{0})D_{\phi,0}(\omega,k)\,$, (33)
[…]
The particle-hole loop calculated with the Green’s functions (31) can be expressed through the Lindhard’s function $\Phi$,
$\intop_{-\infty}^{+\infty}\frac{{\rm d}\epsilon}{2\pi i}\intop\frac{2{\rm d}^{3}p}{(2\pi)^{3}}G_{+}G_{-}=-a^{2}N\Phi(s,x)\,,\,\,s=\frac{\omega}{kv_{\rm F}}\,,\,\,x=\frac{k}{p_{\rm F}}.$ (34)
In the limit $x\ll 1$ the Lindhard’s function is
$\Phi(s,x)\approx 1+\frac{s}{2}\log\frac{s-1}{s+1}-\frac{x^{2}}{12\left(s^{2}-1\right)^{2}}\,.$ (35)
It attains an imaginary part for $s<1$, and in the limit $s\ll 1$ we have
$\Phi(s,x)\approx 1+i\frac{\pi}{2}s-s^{2}-\frac{x^{2}}{12}\,.$ (36)
The spectrum of excitations in the scalar channel, $\omega(k)$, is determined from the pole of $T_{\rm ph,0}$,
$f_{0}^{-1}=-\Phi(s,x)\,.$ (37)
For $1<f_{0}<0$, this equation has only damped solutions with $\Re s<1$ and $\Im s<0$. In the limit of $s,x\ll 1$, using the expansion of (36) we find for the low-lying mode
$\omega_{\rm d}(k)\approx-i\frac{2}{\pi}kv_{\rm F}\frac{1-|f_{0}|}{|f_{0}|}\,$, (38)
which is valid for $1-|f_{0}|\ll 1$ and $k\ll p_{\rm F}$. If $f_{0}<-1$, i.e. the compressibility of the system becomes negative according to Eq. (32), one has $\Im\omega_{\rm d}(k)>0$, and the scalar excitation modes grow exponentially with time as $\propto\exp(+\Im\omega_{\rm d}t)$. This instability is called the Pomeranchuk instability of a Fermi liquid. It may result in the spinodal instability which appears in systems with a van der Waals-like equations of state. Another possibility is the formation of the Bose condensate of a scalar field, in analogy with the Lee-Wick state, but in dilute matter Kolomeitsev and Voskresensky (2016).
We assume that growing unstable bosonic modes may form the Bose condensate described by a scalar boson complex field $\phi$. Such a scalar field can be introduced by means of the Hubbard–Stratonovich transformation or by a formal replacement of the contact interaction with an exchange of an artificial heavy scalar boson Kolomeitsev and Voskresensky (2016). The effective Lagrangian density of the condensate field, taken in the simplest form $\phi=\phi_{0}e^{-i\omega_{c}t+i\vec{k}_{0}\vec{r}}$, can be presented as
$\mathcal{L}_{\phi}=-{\rm sgn}(f_{0})D_{\phi,0}^{-1}(\omega_{c},k_{0})|\phi_{0}|^{2}-{\textstyle\frac{1}{2}}\Lambda(\omega_{c},k_{0})|\phi_{0}|^{4}\,$$\quad$ (39)
$D_{\phi,0}^{-1}(\omega,k)=(\Gamma_{0}^{\omega})^{-1}+a^{2}N\Phi\Big(\frac{\omega}{kv_{\rm F}},\frac{k}{p_{\rm F}}\Big)\,.$
The self-interaction term $\Lambda$ is determined in the leading order by the loop integral of four fermion Green’s functions
[…]
It was evaluated in Brovman and Kagan (1972); Brovman and Kholas (1974) and Dyugaev (1982) in case of one-component fermion system. Energetically preferable is the state with $\omega_{c}=0$ as argued in Ref. Kolomeitsev and Voskresensky (2016). Then we can estimate for $k_{0}\ll p_{\rm F}$,
$\Lambda(0,k_{0})\approx a^{4}\lambda\,\big(1+k_{0}^{2}/2p_{\rm F}^{2}\big)\,,\quad\lambda=\xi\frac{\nu m_{N}^{*\,3}}{\pi^{2}p_{\rm F}^{3}}\,.$ (41)
The factor $\xi$ is introduced to demonstrate uncertainty in the value of the self-interaction parameter $\lambda$, which may arise from other diagrams beyond the leading one in (40). Below we vary it to illustrate sensitivity of our numerical results.
The equation for the condensate amplitude following from the Lagrangian (39) reads
$a^{2}N\widetilde{\omega}^{2}_{s}(k_{0})\phi_{0}+\Lambda(0,k_{0})\,|\phi_{0}|^{2}\phi_{0}=0\,,\quad\widetilde{\omega}^{2}_{s}(k_{0})=|f_{0}(k_{0})|^{-1}-\Re\Phi(0,k_{0})\,.$ (42)
The quantity $|\widetilde{\omega}_{s}(k_{0})|$ plays the role of an effective gap in the excitation spectrum of the scalar quanta similarly to that we considered above in the case of the pion-nucleon interaction. The amplitude of the condensate field, $|\phi_{0}|$, and the Bose condensate energy-density term, $E_{\rm B}$, become for $\widetilde{\omega}^{2}_{s}(k_{0})<0$,
$|\phi_{0}|^{2}=-\frac{N\widetilde{\omega}^{2}_{s}(k_{0})}{a^{2}\lambda}\,\Big(1+\frac{k_{0}^{2}}{2p_{\rm F}^{2}}\Big)^{-1}\,,\quad E_{\rm B}=-\frac{N^{2}\widetilde{\omega}^{4}_{s}(k_{0})}{2\lambda}\Big(1+\frac{k_{0}^{2}}{2p_{\rm F}^{2}}\Big)^{-1}.$ (43)
Simplifying consideration we further put $k_{0}=0$. The case $k_{0}\neq 0$ should be considered separately.
The energy of the system per nucleon including the nucleon part and the contribution from the Bose condensate is as follows
$\mathcal{E}_{N}+\mathcal{E}_{B}=(E_{N}+E_{\rm B}[f_{0}])/n\,.$ (44)
Reference Kolomeitsev and Voskresensky (2016) employed a simple phenomenological parametrization for the volume part of the energy per nucleon, $\mathcal{E}_{N\rm}$ as an expansion in powers of $p_{\rm F}/m_{N}$ proposed in Ref. Kaiser et al. (2002),
$\mathcal{E}_{N\rm}(n)=\frac{3p_{\rm F}^{2}}{10\,m_{N}}-c_{1}\frac{p_{\rm F}^{3}}{m_{N}^{2}}+c_{2}\frac{p_{\rm F}^{4}}{m_{N}^{3}}+c_{3}\frac{p_{\rm F}^{5}}{m_{N}^{4}}\,$, (45)
where parameters $c_{1}$, $c_{2}$ and $c_{3}$ are expressed through values of the binding energy $\mathcal{E}_{0}=-\mathcal{E}_{N\rm}(n_{0})$, compressibility modulus $K$ of the nuclear matter and the nuclear saturation density $n_{0}$. Taking $m_{N}=939$ MeV for the free nucleon mass and appropriate values $n_{0}=0.17\,{\rm fm}^{-3}$, $\mathcal{E}_{0}=-16$ MeV and $K=285$ MeV we find Kolomeitsev and Voskresensky (2016), $c_{1}=3.946$ , $c_{2}=3.837$$c_{3}=13.10$ . From Eq. (45) the volume part of the nucleon contribution to the scalar Landau-Migdal parameter $f_{0}^{N}\equiv f_{0}$ follows as
$f_{0}(n)=\frac{3m_{N}^{*}}{p_{\rm F}^{2}}n\frac{{\rm d}^{2}(n\mathcal{E}_{N,\rm v})}{{\rm d}n^{2}}-1=\frac{m_{N}^{*}}{m_{N}}-1+\frac{m_{N}^{*}}{m_{N}}\frac{p_{\rm F}}{m_{N}}\Big(-6c_{1}+c_{2}\frac{28p_{\rm F}}{3m_{N}}+c_{3}\frac{40p_{\rm F}^{2}}{3m_{N}^{2}}\Big)\,.$ (46)
The energy per particle and the scalar Landau-Migdal parameter are shown in Fig. 8 as functions of the nucleon density. The nucleon volume part of the energy per particle and the parameter $f_{0}$ given by Eq. (46) are presented on the left panel in Fig. 8 by the dotted lines. There is a broad interval of densities, $n_{c}^{\rm(l)}<n<n_{c}^{\rm(u)}$ where $f_{0}<-1$ with $n_{c}^{\rm(l)}=0.422\times 10^{-2}\,n_{0}$ and $n_{c}^{\rm(u)}=0.655\,n_{0}$. The condensate contribution proves to be strongly attractive creating a minimum at $n\sim 0.4\,n_{0}$ for a reduced value of $\lambda$ for $\xi<1/2$. The energies per nucleon $\mathcal{E}_{N}+\mathcal{E}_{B}$ are shown on the left upper panel by thin lines labeled by values of the parameter $\xi$. They are calculated according to Eq. (44) with $f_{0}(n)$ given by Eq. (46).
In the presence of the condensate the incompressibility of the system changes as $K=K_{N}+K_{\rm B}$, $K_{\rm B}=n\frac{d^{2}E_{\rm B}}{dn^{2}}$, and, therefore, the scalar Landau parameter changes too
$f_{0}\to f_{0}^{\rm tot}=f_{0}+\frac{3}{2\epsilon_{\rm F}}K_{\rm B}[f_{0}^{\rm tot}]\,.$ (47)
We used the variational approach to find the solution of Eq. (47), see Ref. Kolomeitsev and Voskresensky (2016), and we fixed the values of $f_{0}^{\rm tot}$ at the ends of the instability interval $f_{0}^{\rm tot}(n_{c}^{(l)})=f_{0}^{\rm tot}(n_{c}^{(u)})=-1$. Thereby we postulated that the condensate arises by the second-order phase transition. Taking into account the possibility of a first-order phase transition, the energy could be further decreased. The energy per particle in our case is constructed using Eq. (32):
$\mathcal{E}_{\rm tot}^{\rm(MF)}(n)=\frac{1}{n}\intop_{n_{c}^{\rm(l)}}^{n}{\rm d}n^{\prime}\intop_{n_{c}^{\rm(l)}}^{n^{\prime}}{\rm d}n^{\prime\prime}\frac{2\epsilon_{\rm F}(n^{\prime\prime})}{3n^{\prime\prime}}(f_{0}^{\rm tot}(n^{\prime\prime})+1)+C\,.$ (48)
The constant $C$ is fixed by requiring $\mathcal{E}_{\rm tot}^{\rm(MF)}(n_{c}^{\rm(u)})=\mathcal{E}_{N}(n_{c}^{\rm(u)})$. The energy (48) corresponds to a mean-field approximation when the reconstruction of the excitation spectrum due to an interaction with the condensate field is not taken into account.
The result of calculations using (48), (47) is shown on the left panel in Fig. 8 by tick lines for different value of the $\xi$ parameters. The self-consitency requirement changes the condensate contribution essentially. The mean-field energy $\mathcal{E}_{\rm tot}^{\rm(MF)}$ becomes now only a smoothly decreasing function of the density and no minimum appears. Nevertheless, as it was expected, the decrease of the $\lambda$ parameter (smaller $\xi$) leads to the decrease of the energy. The density dependence of $f_{0}^{\rm tot}$ is flatter than that of $f_{0}$ for densities $0.2<n<n_{c}^{\rm(u)}$ and $f_{0}^{\rm tot}>f_{0}$. At smaller densities $n\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}0.1\,n_{0}$ we have oppositely $f_{0}^{\rm tot}<f_{0}$. We have to notice that our solutions are accurate at the level of $0.1\%$ for $0.01\,n_{0}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}n\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}0.55\,n_{0}$ but the accuracy is getting worse when density approaches $n_{c}^{\rm(u)}$. Then our solution can be viewed as a smooth extrapolation to $n_{c}^{\rm(u)}$.
Figure 8: Energy per particle (upper panel) and the scalar Landau-Migdal parameter (lower panel) as functions of the nucleon density.
The interaction of the condensate with over-condensate excitations leads to appearance of a new term in the excitation propagator,
$D^{-1}_{\phi,0}\longrightarrow D^{-1}_{\phi}(\omega,k)=D^{-1}_{\phi,0}(\omega,k)-\delta\Sigma_{\phi}(\omega,k)\,$,
where the additional term is
[…]
In such a way we include both mean-field and quadratic fluctuation contributions.
The new spectrum of excitations on top of the condensate is determined now by the equation $D^{-1}_{\phi}(\omega,k)=0$. For $k_{0}=0$ and $1\ll\tilde{\omega}^{2}_{s}(0)<0$ we find
$\omega_{\rm d}\approx i\frac{2}{\pi}kv_{\rm F}\tilde{\omega}^{2}_{s}(0)=i\frac{2}{\pi}kv_{\rm F}\frac{1-|f^{\rm tot}_{0}|}{|f_{0}^{\rm tot}|}\,$,
where we used Eqs. (43) and (41). We pay attention to opposite sign compared to Eq. (38) for $|f_{0}^{\rm tot}|>1$. Thus, we have now $\Im\omega_{\rm d}<0$. Hence, in the presence of the condensate, excitations are damped and the Fermi liquid becomes free from the Pomeranchuk instability.
The particle-hole interaction is also changed in the presence of the condensate. There appears a new term in equation for the particle-hole amplitude (33) after the replacement $D^{-1}_{\phi,0}\longrightarrow D^{-1}_{\phi}$. The new term can be included in the renormalized local interaction
$\frac{1}{f_{\rm 0,ren}}=\frac{1}{f_{0}^{\rm tot}}+2\,\tilde{\omega}^{2}_{s}(0)=-\frac{f_{0}^{\rm tot}}{1+2f_{0}^{\rm tot}}\,$, (50)
where we used $\tilde{\omega}^{2}_{s}(0)=-1-1/f_{0}^{\rm tot}$. Thus, if originally $f_{0}<-1$ and therefore $f_{0}^{\rm tot}<-1$, the renormalized interaction yields $1<f_{\rm 0,ren}<-1/2$. Hence, the first-sound modes becomes stable in the Fermi liquid with the condensate. The quantity $f_{\rm 0,ren}$ is plotted on the right lower panel in Fig. 8. Knowing the value $f_{\rm 0,ren}$ one can reconstruct the energy per particle $\mathcal{E}_{\rm tot}(n)$ of the system using Eq. (48) after the replacement $f_{0}^{\rm tot}\to f_{\rm 0,ren}$ in the integrand. The results are shown on the right upper panel in Fig. 8. We observe the appearance a minimum of the function $\mathcal{E}_{\rm tot}(n)$. For $\xi=1$ the minimum is at $n_{\rm min}=0.085\,n_{0}$ with $\mathcal{E}_{\rm min}=-16.01$ MeV. Variation of the parameter $\xi$ in interval $\xi=1/4\mbox{--}2.0$ leads to variations of the values $n_{\rm min}=(0.82\mbox{--}0.93)\times 10^{-1}\,n_{0}$ with $\mathcal{E}_{\rm min}=-(16.8\mbox{--}15.6)$ MeV. So the minimum in energy can correspond to a new stable or metastable state of the dilute nuclear matter compared with the usual saturation energy of 16-16 MeV used to specify the equation of state of the ISM.
Finally, we should mention that the obtained results are valid, if on the one hand $\phi_{0}$ is rather small and on the other hand fluctuations on the top of the condensate yield a yet smaller contribution, since the self-interaction of excitations on top of the condensate and feedback of fluctuations on the mean field were disregarded.
Considering finite system one must take into account the surface and Coulomb energies. We parameterize the surface energy as in Migdal et al. (1976); Migdal (1978),
$\mathcal{E}_{\rm S}A=a_{\rm S}A^{2/3}\,,\quad a_{\rm S}\approx C_{\rm S}\Big(\frac{n}{n_{0}}\Big)^{1/3}\frac{\mathcal{E}(n,z)}{\mathcal{E}(n_{0},z=1/2)}\,,\quad C_{\rm S}=18\,{\rm MeV}\,$, (51)
where $z=Z/A$.
With the account for the surface energy (51) the Landau-Migdal parameter, $f_{0}^{\rm S}$, takes the following form
$f_{0}^{\rm S}(n)$ $=f_{0}(n)+\frac{C_{\rm S}}{A^{1/3}}\frac{1}{\mathcal{E}_{N,\rm v}(n_{0})}\Big(\frac{n}{n_{0}}\Big)^{1/3}\frac{m_{N}^{*}}{m_{N}}\Big(\frac{3}{5}-c_{1}\frac{28p_{\rm F}}{9m_{N}}+c_{2}\frac{40p_{\rm F}^{2}}{9m_{N}^{2}}+c_{3}\frac{6p_{\rm F}^{3}}{m_{N}^{3}}\Big)\,.$ (52)
Following  Kolomeitsev and Voskresensky (2016), as an example, we consider a system of 125 nucleons at $N\approx Z$. Then the energy per particle $\mathcal{E}_{N}+\mathcal{E}_{\rm S}$ saturates at density $0.95\,n_{0}$ with the minimum $\mathcal{E}_{0}^{\rm S}=-12.4$ MeV. With taking into account the surface energy and solving Eq. (47) with $f_{0}^{\rm S}$ instead of $f_{0}$ we find that the new minimum appears at $n_{\rm min}=0.74\times 10^{-1}\,n_{0}$ with the depth $\mathcal{E}_{\rm min}=-13.04$ MeV. So, $\mathcal{E}_{\rm min}<\mathcal{E}_{0}^{\rm S}$ and, hence, a new stable state of matter is possible in the finite size object with account for the surface energy. To analyze its stability against fusion we must include the Coulomb energy
$\mathcal{E}_{\rm C}A\simeq a_{\rm C}z^{2}A^{5/3}\,,\quad a_{\rm C}=C_{\rm C}\Big(\frac{n}{n_{0}}\Big)^{1/3}\,,\quad C_{\rm C}=0.7\,{\rm MeV}\,.$ (53)
A spherical nucleus is stable against an ellipsoidal deformation (stretching) if
$\frac{Z^{2}}{A}<2\frac{a_{\rm S}(n_{\rm min})}{a_{\rm C}(n_{\rm min})}=2\frac{C_{\rm S}}{C_{\rm C}}\frac{\mathcal{E}_{\rm min}}{\mathcal{E}_{0}}\simeq 42\,.$ (54)
We see that, because of the condensate formation, a new metastable dilute nuclear state is stable against spontaneous the fission for $A\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}170$.
Reference Maslov and Voskresensky (2019) studied the equation of state of isospin-asymmetric matter and demonstrated that even for the asymmetry parameter $n_{p}/n_{n}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}0.2$ there remains a broad region of proton and neutron densities where $f_{0}<-1$ and, therefore, the spinodal instability may develop. For a nugget of a size $R$ much larger than the Debye screening length $l_{D}\simeq\sqrt{\pi}/(2|e|V_{0})$, the Coulomb contribution from protons must be screened by electrons, $n_{p}=n_{e}$, cf. Migdal et al. (1977). Here we for simplicity assume that the scalar field is electrically neutral. The electron Fermi momentum is then $p_{{\rm F},e}=V_{0}=(3\pi^{2}n_{p})^{1/3}$ and the electron kinetic energy density is $E_{e}=p_{{\rm F},e}^{4}/(4\pi^{2})$. In the isospin-asymmetric matter, the nuclear symmetry energy must be included. A simple parametrization was proposed in Heiselberg and Hjorth-Jensen (1999), which is a fit to microscopic calculations
$E_{N,{\rm sym}}(n_{p},n_{n})=J\,\frac{(n_{p}+n_{n})^{\gamma}}{n_{0}^{\gamma}}\frac{(n_{p}-n_{n})^{2}}{n_{p}+n_{n}}\,,\quad J=32\,{\rm MeV}\,,\quad\gamma=0.6\,.$ (55)
Let $n_{p}=\chi n_{n}$, then the total energy per particle of the system is
$\mathcal{E}(\chi)$ $=\mathcal{E}_{\rm min}+J\,\Big(\frac{n_{\rm min}}{n_{0}}\Big)^{\gamma}\frac{(1-\chi)^{2}}{(1+\chi)^{2}}+\frac{3}{4}(3\pi^{2}n_{\rm min})^{1/3}\frac{\chi^{4/3}}{(1+\chi)^{4/3}}$
$\approx\mathcal{E}_{\rm min}+7.3\,{\rm MeV}\frac{(1-\chi)^{2}}{(1+\chi)^{2}}+149\,{\rm MeV}\frac{\chi^{4/3}}{(1+\chi)^{4/3}}\,$, (56)
where $n_{\rm min}=0.85\times 10^{-1}\,n_{0}$ is used. If we take $\chi\sim 0.2$ we obtain $\mathcal{E}(\chi\sim 0.2)\approx\mathcal{E}_{\rm min}+16.9\,{\rm MeV}$. Minimizing $\mathcal{E}(\chi)$ with respect to $\chi$ we find that the minimum is realized for $\chi_{\rm min}\approx 3.1\times 10^{-3}$ and in the minimum $\mathcal{E}(\chi_{\rm min})\approx\mathcal{E}_{\rm min}+7.3\,{\rm MeV}$. Thus if $\mathcal{E}_{\rm min}$ is of the order (1017)-(10-17) MeV in the case of the isospin asymmetric nuclear matter, one may rise the question about a possibility of the existence of neutron-rich dilute compact stars glued by an collective scalar field. However, the above estimates crucially depend on the value of symmetry energy of nuclear matter at small densities: the smaller $E_{N,{\rm sym}}$ is, the smaller will be repulsive contributions in Eq. (56).
Above we only demonstrated a possibility of the scalar field condensation in the dilute nuclear systems, which may result in occurrence of new metastable or stable states. A more detailed analysis is still required to justify this hypothesis.
7.2 Bose condensates of nuclear clusters
At densities when $f_{0}(n)<-1$ the incompressibility of nuclear matter is negative and, as we have mentioned, not only to the scalar particle-hole mode condensation but also the spinodal instability of matter may occur. In the low energy heavy-ion collisions this instability manifests itself in an abrupt increase of the production of $\alpha$ particles and other nuclear clusters, see Röpke et al. (1982); Curtin et al. (1983); Bertsch and Siemens (1983); Schulz et al. (1983); Maslov and Voskresensky (2019); Borderie et al. (2022). With a binding energy of 7.07 MeV per nucleon $\alpha$ particles are bound as strongly as the most tightly bound nucleus of 52Fe. Under certain conditions $\alpha$ particles can form a stable Bose-Einstein condensate state (see Ref. Typel et al. (2010); Schuck et al. (2016) and references therein). Reference Clark and Krotscheck (2024) argued that the spectrum of excitations in the system of interacting $\alpha$ particles is similar to the spectrum of He4 and has the roton minimum at $k\neq 0$. If so, the inhomogeneous $\alpha$ condensate occupying the roton minimum could occur in the moving (e.g. rotating) pieces of the $\alpha$ particle system. Similar effects were studied in Refs. Voskresensky (1993, 2023); Kolomeitsev and Voskresensky (2017). Also, one considers the possibility of deuteron condensation in dilute nuclear systems as discussed in Ref. Lombardo et al. (2001), despite the fact that deuterons are seven times weakly bound than $\alpha$ particles.
Consider an extended nucleon system of large size and study whether it is possible to form a Bose-Einstein condensate from bosonic clusters consisting of an even number of nucleons. We must take into account an electric field and $\beta$ equilibrium conditions, as it was done in Ref. Voskresensky (1977); Migdal et al. (1990) and discussed above in Section 3.
Let $A$ is the total number of nucleons in a piece of dilute nuclear matter containing the Bose-Einstein condensate of clusters consisting of $A_{\rm cl}$ nucleons each (e.g. $A_{\rm cl}=4$ for the $\alpha$ particle). The cluster condensate is described by a complex field $\phi$. The Lagrangian density of a stationary, homogeneous condensate field $\phi$ can be written as
$\mathcal{L}_{\rm cl}=(\omega+V)^{2}|\phi|^{2}-m^{2}_{\rm cl}|\phi|^{2}+n_{e}V-\frac{V^{4}}{4\pi^{2}}\,$, (57)
where $\omega+V$ is the energy of a positively charged cluster quasiparticle shifted by the electric potential $V$. Considering a large system, we may drop gradient terms with $\nabla\phi$ and $\nabla V$ and neglect for $\phi$ self-interaction terms for simplicity. The equation of motion for the scalar field following from the Lagrangian (57) is $(\omega+V)^{2}\phi=m^{2}_{cl}\phi$. The condition of the charge neutrality of the system requires that the positive charge of clusters is compensated by the charge of electrons, therefore $n_{\rm cl}=n_{e}=V^{3}/(3\pi^{2})$, where the number density of cluster quasiparticles is
$n_{\rm cl}=\frac{\partial\mathcal{L}_{\rm cl}}{\partial\omega}=-2(\omega+V)|\phi|^{2}\,.$
The energy density of a cluster is $A_{\rm cl}\big(m_{N}-\mathcal{E}_{\rm cl}^{\rm(bind)}\big)n_{\rm cl}$, where $\mathcal{E}^{\rm(bind)}_{\rm cl}>0$ is the cluster binding energy per nucleon, e.g., $\mathcal{E}_{\rm cl}^{\rm(bind)}\simeq 7.07$ MeV for the $\alpha$ particle. Then the energy density of the condensate of the bosonic clusters can be presented as
$\Delta E=E_{\rm tot}-A_{\rm cl}m_{N}n_{\rm cl}\simeq\frac{V^{4}}{12\pi^{2}}-A_{\rm cl}\mathcal{E}_{\rm cl}^{\rm(bind)}n_{\rm cl}=\Big(\frac{1}{4}(3\pi^{2}n_{\rm cl})^{1/3}-A_{\rm cl}\mathcal{E}_{\rm cl}^{\rm(bind)}\Big)n_{\rm cl}\,.$ (58)
Thus, in this simplified model, the system is unbound ($\Delta E>0$) at densities,
$n\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}(4A_{\rm cl}\mathcal{E}_{\rm cl}^{\rm(bind)})^{3}/(3\pi^{2})$
in the large system limit. For $\alpha$ particles it occurs at densities $n>0.04\,n_{0}$.
7.3 Dilute nonequilibrium states
The production of pion-enriched matter occurs in ultrarelativistic heavy-ion collisions at highest RHIC and LHC energies. At expansion of the system, the temperature of the pion gas decreases. At a certain temperature, $T_{chem}$, the particle populations stop changing, reaching a state called chemical freeze-out, though the particles still remain in the thermal equilibrium. Then, at a later stage of the expansion up to a thermal freeze-out, till $T(t)>T_{th}$, we are dealing with a system with a dynamically fixed number of pions. At $T<T_{BEC}$, provided $T_{th}<T_{BEC}$, there may appear the Bose-Einstein condensate of pions characterized by the dynamically fixed particle number  Voskresensky (1994); Kolomeitsev et al. (2021).
Another interesting situation is possible: in peripheral heavy-ion collisions, the colliding nuclei can be considered as almost freely interpenetrating beams of nucleons. At the collision energy $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}300$ MeV per nucleon, the nucleon Fermi spheres from each beam do not overlap in momentum space. Since the pion polarization operator is $\Pi\propto-\nu p_{\rm F}k^{2}$, as discussed in Section 5, this corresponds to a density which is effectively four times higher than in the case of overlapping Fermi spheres. Thus, there is a possibility of observing the effects of the $p$ wave pion condensation in peripheral heavy-ion collisions Pirner and Voskresensky (1995); Voskresensky (2025).
7.4 Dilute quantum droplets
The phenomenon of Efimov states in Fermi systems opens another interesting possibility: For particular values of the diluteness parameter $n|a|^{3}$, where $n$ is the fermion density, $a$ is the scattering length, a three-body interaction can provide a dominant contribution in the energy-density functional Bulgac (2002). if this happens, both bosonic and fermionic systems of various sizes and densities could become self-bound. In this relation, we note that the nucleon-nucleon scattering length has an anomalously large magnitude, $a_{NN}\simeq-20$ fm. A discussion of the role of this circumstance for description of a dilute neutron matter can be found in Kolomeitsev et al. (2024). Finally, we notice that some experiments claimed an observation of a tetraneutron ($n^{4}$) signal, which may be associated with a bound state or an unbound resonance state, see discussion in Faestermann and Gernhäuser (2025).
As we have mentioned, in the spectrum of $\pi^{+}$ in neutron matter in the interval of densities $n_{c}^{\pi^{\pm}}>n>n_{c}^{\pi^{+}}$, where $n_{c}^{\pi^{+}}<n_{0}$ according estimate of Migdal et al. (1974), and $n_{c}^{\pi^{\pm}}\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}n_{0}$, there appears the branch with $\omega^{\pi^{+}}(k)<0$ at $k\neq 0$. Appearance of such a branch may result in some important consequences, see Voskresensky (1993). For instance, at the formation of the neutron star in a supernova explosion a part of the initial angular momentum can be absorbed by the $\pi^{+}$ condensate subsystem. Metastable $\pi^{+}$-multi-neutron droplets with a large angular momentum and a density $n>n_{c}^{\pi^{+}}$ can be formed. They may take away part of the total angular momentum of the star. Metastable $\pi^{+}$ condensate neutron-rich droplets also can be formed with some probability during heavy-ion collisions. Since density in such droplets might be smaller than $n_{0}$ they may behave as anomalons mentioned above, see Voskresensky (1993). Also there a possibility of appearance of condensate fragments with a large angular momentum in peripheral heavy-ion collisions was discussed.
8 Observations
Search of abnormal nuclei, nuclearites and nuclei-stars composed of absolutely stable matter is an intriguing prospect in astrophysics. Detection of such objects is not a trivial problem due to numerous observational signatures, which are superimposed on those of ordinary neutron stars and white dwarfs. Distinguishing features may be found in ultra-fast rotation (well below 1 ms), star radii less than 9-10 km, specific photon emission from a star surface, etc. Possible candidates among compact objects were identified: RX J1856.5-3754, PSR J1614-2230, PSR J0348+0432, 4U 1820-30, HESS J1731-347, XTE J1814-138. However, definitive conclusions whether these objects are anomalous or not, could not be drawn. Additionally, particular measurements of the low-mass and small-radius compact star HESS J1731-347 depend heavily on the assumed atmosphere model and are sensitive to errors in distance and inferred age. Let us add here the latest results of NICER Mauviard et al. (2025) for PSR J0614-3329 giving the rather small equatorial radius $10.29^{+1.01}_{-0.86}$ km for the mass $M=1.44^{+0.06}_{-0.07}\,M_{\odot}$ (median values with equal-tailed $68\%$ credible interval).
Experimental search for abnormal nuclei was carried out in natural samples among the products of interactions between high-energy particles and matter, in fission products, in heavy-ion collisions, etc., see Refs. Frankel et al. (1976); Holt et al. (1976); Aleshin et al. (1976); Kulikov and Pontecorvo (1977); Price and Stevenson (1975); Frankel and Stevenson (1976); Karnaukhov and Polikarpov (1977); Buchwald et al. (1981). Although no abnormal states were found presently, a number of observations demonstrate anomalies that defy conventional physics interpretations. May be at least some of them are related to the anomalous nuclear states. We mention several such anomalies. A very heavy particle was detected passing through a balloon-borne stack of Cherenkov film, emulsion, and Lexan sheets Price et al. (1975). It could be a neutron anomalous nucleus. The spectra of nuclei obtained in the aerostatic JACEE experiment with an emulsion chamber show a peculiar bend at the cosmic ray energy of 3 PeV. The change in the slope of the spectrum was suggested in Ref. Shaulov et al. (2022) to be explained by the appearance of an anomalous component of cosmic rays, which is much heavier than the nuclear component and, possibly, consists of particles of strange quark matter. In the transition region, both nuclei and strangelets can exist, that leads to large fluctuations in the events observed in the experiments with X-ray emulsion chambers.
Recent works Bertolucci et al. (2017); Zioutas et al. (2020, 2023) argued that there are other observations, which can be hardly explained in the framework of conventional physics. These anomalies include: unexpected correlations with temperature variations in stratosphere, the total electron content of the Earth atmosphere, correlations between earthquake activity and relative positions of planets, unexpected seasonal stratospheric temperature variations, ionospheric anomalies, and their correlations with seismic activity, as well as solar phenomena like the coronal heating problem, the origin of sunspots, and the trigger mechanism of the solar flares, see also Maroudas (2022); Zioutas et al. (2022). The statistical significance of the observed correlations is high. Reference Zhitnitsky and Maroudas (2025) suggested that in mentioned events one may deal with hypothetical very dense and microscopically large composite objects with the mass of the order of grams and sub-micrometer size consisting of nuggets of strongly interacting matter. For instance, it was suggested to look for so-called solar nano-flares with the axion quark nugget annihilation events in the solar corona. Note that, if instead of quark nuggets one considered the Lee-Wick anomalous state or the Migdal pion condensate state abnormal nuclei, at least a part physical consequences would be similar to those for strangelets.
Let us now mention another possibilities. In a uniform electric field with the strength of $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10^{4}$ V/cm the depth of the potential $V_{0}$ exceeds $2m_{e}\approx 10^{6}$ eV when the distance between plates is $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10$ cm. For a 300  times larger distance between plates $V_{0}$ exceeds two pion masses, $2m_{\pi}\approx 280$ MeV. For comparison, the breakdown voltage for the dry air at atmospheric pressure is $(2\mbox{--}3)\times 10^{4}$ V/cm and for some types of glasses it reaches $3\times 10^{6}$ V/cm. Similar estimates can be found for the cylindrical and spherical capacitors. However we must stress that despite the potential inside the capacitor may readily exceed $2m_{e}$ and $2m_{\pi}$ respectively, the probability of the spontaneous pair production from the vacuum is exponentially small in both cases, $\propto e^{-m^{2}/E}$. Nevertheless, electrons and charged pions can be produced near the walls of the capacitor with a larger probability provided the electric potential well is sufficiently deep. In the broad potential well for $\pi^{-}$ in the region, where $V>m_{\pi}$, the $\pi^{-}$ ground state energy level crosses $\omega=0$ and formation of the static $\pi^{-}$ condensate field becomes energetically favorable. In the region with $V>m_{\pi}$ the static $\pi^{+}$ condensate field can be formed. Note that in the presence of a charged condensate, photons become effectively massive particles via the Anderson-Higgs mechanism. Therefore, the condensate region should look dark. This circumstance can help in identifying the effect.
It is worth noticing that deep potential wells can be formed in thunderstorms on Earth. During tens of minutes before a discharge, the electric voltage may reach values of $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10^{8}$ V at distance $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}(1-10)$ km between clouds, cf. Marshall and Stolzenburg (2001). Hence the typical potential well reaches values $V_{0}\gg m_{e}$ at such distances. An excess of neutrons during thunderstorms has been indeed detected, however it was associated with photonuclear reactions Babich (2019). Strong thunderstorms may also occur in atmospheres of other planets such as Jupiter and Venus.
Finally, let us notice that 10 anomalous bursts correlated with lightnings have been observed in thunderstorms in 5 years Abbasi et al. (2017); Okuda (2019). The estimated energy from individual events within the bursts is five to six orders of magnitude higher than the energy estimated by an event rate. The multiple air showers occurred within 1 ms, defying explanation by conventional high-energy cosmic rays. Similar “exotic events” were recorded by the AUGER collaboration. The ANITA experiment detected two anomalous upward-propagating events with non-inverted polarity. The list of anomalous atmospheric events also includes the Multi-Modal Clustering Events observed by HORIZON 10T.
If a nugget of a new phase of matter passes through the thunderstorm clouds with an ultra-high voltage ($\sim 10^{9}$ V) it will be accelerated and its electron shell will be destroyed in possible collisions with a release of a significant energy and baryon production.
9 Concluding remarks
In 1971 Migdal suggested possibility of existence of metastable or stable abnormal pion condensate nuclei and Bodmer assumed possibility of collapsed quark nuclei. In 1974 Lee and Wick proposed scalar (sigma) condensate abnormal nuclei. In 1984 Witten suggested quark nuggets. These pioneering suggestions paved the way for numerous subsequent theoretical and experimental investigations of various anomalous nuclear objects. New more detailed models were constructed. Nevertheless, in spite of more than fifty years were past, as in early days of this field, there remain essential uncertainties in theoretical predictions due to a lack of our knowledge of a strongly interacting dense matter. In spite of abnormal nuclear states were not observed, the number of observations grows, which are hardly explained by conventional physics. Some of such events could be associated with manifestation of anomalous nuclear systems. Let us be patient. Time flows not in vain.
[…]
All authors have equally contributed to the research and preparation of the manuscript.
[…]
EEK acknowledges the support by the Slovak grant VEGA 1/0521/22 and the support within the framework of the scientific program of the National Center for Physics and Mathematics in Sarov, topic No. 6 "Nuclear and Radiation Physics" (stage 2023-2025).
[…]
Not applicable
[…]
Not applicable
[…]
No explicit data set are prepared.
[…]
The authors declare no conflict of interest.
[…]
Abbreviations The following abbreviations are used in this manuscript:
EoS Equation of State
HESS High Energy Stereoscopic System
ISM Isospin-Symmetric Matter
NICER Neutron Star Interior Composition Explorer
PCAC Partial Conservation of Axial Current
PNM Pure Neutron Matter
PSR Pulsating Source of Radiation
QCD Quantum Chromo-Dynamics
RMF Relativistic Mean Field
XTE Rossi X-Ray Timing Explorer
4U Fourth UHURU Catalogue
[…]